A Complete Bibliography of Zeitschrift für allgemeine Wissenschaftstheorie / Journal

for General Philosophy of Science

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

05 June 2023 Version 1.04

Title word cross-reference

· [542].

-geschichte [520].

18th [515]. 1919 [332]. 1953 [382]. 1979 [396].

= [108].

A. [540]. abgeschlossene [335]. abgeschlossenen [337]. Abstract [652]. Abstrakte [652]. according [521]. Account [423]. Accumulative [538]. Acham [195]. Achinstein [128]. action [508, 217, 508, 383]. action-theoretical [508]. Acts [256]. ad [612]. Adäquatheit [266]. Addis [235]. Adorno [74, 296]. after [249]. Agassi [657, 600]. Akademie [76]. Aktualität [411]. Alan [73, 500]. Alasdair [108]. Albert [443, 332, 177]. aller [619]. allgemeine [443]. Allgemeinen [566, 19, 650, 550]. alltagsweltlicher [265]. älteren [489]. Alwin [588, 663, 588, 663, 563]. am [26, 571, 285, 475]. America [511]. Analogy [368]. Analyse [185, 349, 526]. Analysemodelle [439]. Analysen [499]. Analyses [94]. Analysis [510, 316, 338, 522, 540, 242, 525, 526, 500, 530, 583]. Analytic [657, 125]. analytical [576, 560, 543, 217]. analytische [346, 560, 543, 195, 105]. Analytischen [42, 10, 392, 386, 412, 243, 121]. analytischer [83, 188]. anarchistischen [304]. Anatomie [542]. anatomy [542]. Anderson [380]. Andersson [481]. Andreas [392]. Anglistik [329]. Anlaß [396, 556]. Anmerkung [392]. Anmerkungen [153, 355, 214, 357]. Anschauung [179]. Anschluß [452]. Ansprüche [318]. anthropologischen [404]. antiken [160]. Antinomie [283]. Antinomienproblem [170]. Antinomienproblems [212]. Antwort [494]. Anwendung [191]. Application [256]. Approach [565]. Approximationstheorie [244]. äquivalente [557]. Arbeit [439]. Arbeiten [76]. arbeitet [346]. Argument [211, 596]. Argumentation [539, 377, 187]. Argumentationsfiguren [415]. Argumentationsspiel [596]. argumentationstheoretischer [415]. argumentative [601]. argumentativen [601]. Argumente [276]. Arguments [299]. Aristotle [131]. arithmetic [576]. Arithmetik [576]. Armin [142]. Aron [130]. Arthur [217]. Artifical [555]. Artificial [605]. Assessment [500, 509]. Atomic [423]. Atomphysics [54]. Atomphysik [54]. Attempt [528, 571]. Audretsch [658]. Auffassung [635]. Aufgabe [619]. Aufgaben [83, 121]. Aufklärung [107]. Auflage [393]. Auflösung [598]. Aufsatz [65]. August [68, 158, 69]. Auseinandersetzung [245, 474, 393]. Aussagen [155, 99]. Aussagenlogik [381]. Ausverkauf [452]. Autonomieanspruchs [2]. Autorität [342]. Axiomatische [285, 387]. Axiome [208]. Azari [379].

B [604, 178, 25]. Back [29, 58, 80, 111, 144, 237, 273, 288, 399, 430, 484, 502, 562, 590, 609, 624, 642, 323]. Background [12]. Backhaus [247]. Bacon [344]. Bamberg [425]. Band [140, 42, 72, 10]. Bar [274]. Bar-Hillel [274]. Basic [59, 360, 602]. Basis [532, 387, 521]. Baum [640]. Baumgartner [303]. Bayertz [659]. Bedeutung [650, 277, 179, 499, 525, 628, 280]. Bedeutungsverstehen [634]. Bedingungen [172]. Bedingungssätze [259]. Bedürfnis [378]. Beeh [477]. Beendigung [186]. Begriff [337, 261, 179, 243]. Begriffe [504, 580, 291, 277, 387]. Begriffen [222]. begriffs [330]. begriffs- [330]. Begründung [245, 63, 503, 548, 559, 264, 412, 472, 488, 10]. Begründungen [206]. Begründungsphilosophischer [313]. being [524, 524]. Beispiel [26, 571, 285, 615]. Beitrag [326, 415, 635, 206, 436, 376, 72]. Beiträge [269]. Beiträgen [155]. Bela [56, 45]. Belgium [181, 176]. Bellsche [257]. Bemerkung [233, 337, 33, 358]. Bemerkungen [245, 190, 258, 635, 601, 150, 473, 89, 246, 412, 156, 65, 291, 442, 212, 580]. Benennung [551]. Benjamin [478]. Benzolformel [84]. Beobachter [340]. Bericht [476, 37, 68, 158, 69]. Berka [497]. Berlin [477]. Berlin/New [477]. Bernard [179]. Bernays [275]. Beschreibung [49]. Besteigung [619]. Bestimmung [149]. betrachtet [471, 533]. Betrachtung [88]. Betrachtungen [559]. Better [463]. between [256, 298, 615, 618, 175, 586, 449]. Bewährung [459, 613, 389]. Bewältigung [619]. Bewegungsaxiom [260]. Beweis [233, 257, 172]. Bewertung [614]. Bewertungen [282]. Bewußtseinsphilosophie [542]. Bibler [25]. Bibliographie [13, 43, 56, 78, 663, 641, 362, 398, 429, 454, 483, 501, 608, 516, 531, 544, 589, 180, 269, 182, 219, 253, 287, 322, 110, 143, 588, 428]. Bibliography [516, 181, 663, 588, 641, 608, 531, 544, 589]. Bildung [222]. Biological [640, 586]. biologie [457]. biologischen [640]. biologischer [452, 477, 586]. biology [457]. Birjukov [178]. bis [69]. Black [372]. bleibt [355]. blocks [601]. Body [434, 604]. Böhler [542]. Böhlers [542]. Böhme [637]. Bohr [465, 423]. Bohrs [465]. Boltzmann [487]. Bolzanos [179]. Book [558, 73, 93, 252, 604, 94, 360, 383, 622, 108, 178, 497, 126, 606, 12, 41, 479, 478, 305, 162, 500, 530, 125, 127, 128, 197, 217, 235, 216, 605, 638, 639, 123, 129, 341, 395, 427, 512, 583, 25, 382, 426, 54]. Books [545, 561, 623]. Bootstrap [96]. Boscovich [519, 519]. Bound [632]. Branch [433]. Braun [361]. breaking [617]. Brennpunktproblemen [124]. Brief [332]. bringt [355]. Britain [159, 496]. Bruno [513, 426]. Bubner [141, 319]. Büchel [413]. Bücher [164, 201, 236, 272, 307, 343, 384, 416, 444, 468, 545, 561, 623]. Buffon [366]. Bühlers [285]. Bukarest [69]. Bundesrepublik [340]. Bunge [123, 129, 341, 395]. Bußhoffs [441]. Butts [360]. bzw [375].

C [217, 382]. C. [521]. called [524]. Can [424, 638, 509]. Canada [157]. Canadian [163]. Canaris [26]. Capacities [279]. Carl [93]. Carnap [30]. Carpenter [500]. Carter [604, 604]. Cartesian [604]. Cäsar [333]. Case [223, 523, 649, 550]. Cassirer [62]. Causality [168, 566, 568]. Causation [242]. Cause [556]. Causes [558]. Century [515]. Ceteris [345]. Chance [559]. Chancengewichtungen [350]. Change [173, 567, 619, 515, 93]. Character [655]. Chemie [515]. Chemistry [433, 572, 515]. Chinese [603]. chinesischer [603]. Chomsky [593]. Christof [251]. citation [525]. civilization [542]. Classical [252]. Classifications [314]. Claus [26]. Claus-Wilhelm [26]. Clauses [345]. Climbing [619]. Closure [345]. Cognition [256, 640]. Cognitive [533, 424, 574, 533]. cognitivism [604]. Cohen [382]. Coherence [162]. Cojectivity [100]. Collective [127]. Color [655]. Comment [600]. Commentary [138, 137]. Comments [138, 174, 635]. Communication [82, 542, 530]. comparison [582]. Complementarity [62]. complete [537]. Completed [440]. Compound [434]. Comprehensive [583]. Computability [360]. Concept [315, 526]. conceptions [567]. Concepts [497, 580, 504, 616, 550, 127]. Concise [9]. conditions [598]. Conflict [566, 636]. Confrontation [493]. Confusions [633]. Congress [360, 603]. Conjectural [424]. Consciousness [542]. Consensus [541]. Conserving [434]. Consideration [655, 576]. construct [552]. Construction [426]. constructive [579]. contemporanea [482]. Contemporary [159, 119, 41]. content [550]. Context [458, 136, 137]. Contraction [495]. contribution [635]. Controversy [298]. Copenhagen [537]. Coreth [72]. Corroborate [308]. Cosmology [478]. Counter [299]. Counter-Arguments [299]. Cramer [141]. Criteria [612]. Criterion [656]. Critical [338, 559, 118, 580, 102, 576]. Criticism [466, 338, 423, 465, 73]. Critique [252, 338, 656, 585, 542]. Cross [535]. Cross-Fertilization [535]. cultura [482]. cultural [586]. Culture [500, 632]. Culture-Bound [632]. Cultures [630]. Cunningham [216]. current [536]. Cycle [540].

D [542, 396]. Danto [217]. Darstellung [392]. David [587]. dazu [116]. de-finalization [578]. Debatte [411]. Deductive [463]. deduktiv [226]. deduktiv-nomologischer [226]. deduktive [3]. Definierbarkeit [259, 33]. Definieren [198]. Definition [353, 491]. Degenerating [423]. degree [613]. Del [52]. Del-Negro [52]. Demarcating [647]. Denken [113, 27]. Denkfiguren [448]. Denkmethoden [443]. Denkstil [313]. Denmark [23]. dennoch [557]. deontischen [120]. Derrida [628, 628]. Descarte [113, 604]. Descartes [434, 46, 60]. Detel [354]. determining [331]. Deutschen [442, 26, 67]. Deutschland [340]. Deutung [294, 537]. Development [249, 312, 317, 25, 567, 619]. Developments [91]. Devil [367]. Dezisionismus [264]. Dialektik [98, 88, 141, 101]. dialektischen [88]. Dialog [520, 344]. dialogical [542]. dialogischen [542]. Dialogue [520]. Diederich [196]. Diemer [563, 588, 663]. Diesseits [331]. Differenz [265]. Dilthey [40, 585]. Dimensions [360]. Disagreement [295]. Discipline [329]. Discourse [530, 317]. Discovery [458, 602, 491]. Discussed [655]. Discussion [612, 450]. Disentangling [299]. Diskurstheorie [171]. Diskussion [241, 214, 358]. Dispositional [593]. Dispositionsbegriffe [35]. Dispositionsprädikaten [259]. dispute [536]. Distance [368]. Distinction [256, 125]. Disziplinen [225]. Diversity [317]. do [509, 638]. does [552]. Dogma [451, 467, 393]. Dogmatismus [85]. Drache [355]. draft [601]. dreißigjährige [38]. Dualismus [294]. Duerr [453]. Durbin [500]. Düsberg [397].

e.V [425]. Eastern [9]. Eberlein [578]. Eberleins [578]. Eckard [370]. ecology [550]. Economic [252]. Economics [248, 252]. Edgar [480]. Editorial [645]. Editors [94]. Edward [252]. Efficient [558]. Eichner [298, 211]. Eigenschaften [224]. Eike [198, 271, 377]. Eindeutigkeit [579, 232]. einem [601, 65]. Einführung [443, 55]. Eingegangene [164, 201, 236, 272, 307, 343, 384, 416, 444, 468, 545, 561, 623]. Einheit [628, 617]. Einige [412, 190, 113, 556]. einigen [499]. Einstein [332]. Einsteins [469, 332]. Einzeldarstellungen [72]. einzelsprachlich [408]. Einzelwissenschaften [285]. Electronics [407]. Elementare [443, 390, 470]. Elementarteilchenphysik [469]. Elementary [223, 523]. Elend [74]. Eliminierung [170]. Elisabeth [515]. Emanzipation [17]. emanzipatorischen [89]. Emerich [72]. Empfehlungen [282]. empirical [552, 581, 654]. empirical-scientific [552]. empirically [557]. Empiricism [566, 125]. empiricist [539, 382]. empirisch [552, 557]. empirisch-wissenschaftliche [552]. Empirische [268, 309, 389, 654, 276]. empirischen [475, 376]. empirischer [488, 581, 185]. Empirismus [566, 117]. empiristischer [539]. Ende [412]. Engelhardt [498]. Entdeckung [153, 84, 47]. Entfinalisierung [578]. Entgegnung [465]. Entropie [568]. entropy [568]. Entscheidung [339]. entscheidungstheoretische [191]. entschleiert [586]. entwickelt [26]. Entwicklung [619]. Entwurf [601, 49, 250]. Epistemological [605, 650, 613, 507]. Epistemology [620, 509, 248]. equivalent [557]. Ereignis [262]. Erfahrung [165, 42, 85]. Erfahrungswidrigkeit [192]. erfindende [113]. Ergänzungen [56, 180, 411]. Ergebnissen [550]. Erkennens [640]. Erkenntnis [379, 650, 436, 414, 396, 640, 321, 339, 342]. Erkenntnisfortschritt [258, 150, 354]. Erkenntniskritik [101]. Erkenntnismodell [613, 199]. Erkenntnisprozessen [436]. erkenntnistheoretische [179]. erkenntnistheoretischen [650]. erkenntnistheoretischer [222]. Erkenntnistheorie [452, 509, 156, 35, 620, 473, 304, 348]. Erklären [209, 661]. Erklärung [629, 533, 10, 22]. Erklärung. [124]. Erklärungen [265, 438, 553]. Erklärungsbegriff [226]. Erklärungsmodelle [614]. Erklärungstheorie [352]. Erläuterung [648]. Erleben [122]. Ernest [292]. erneute [410]. Ernst [567, 177, 567]. erstellt [49]. Erster [603]. Ertrag [284, 302]. Erwiderung [300, 247]. Erziehungswissenschaft [370]. Essay [405, 128, 639, 169, 512]. Essays [480]. Essler [105]. Ethics [559, 312, 618]. Ethik [559, 120]. europäischen [330]. Europe [9]. evaluate [614]. Evaluation [315]. Evolution [567, 127, 586]. Evolutionäre [567, 509, 452]. Evolutionären [452, 620, 473]. Evolutionary [620, 506, 509, 567]. Exact [108]. Exaktheit [297]. example [571, 615]. exoterische [326]. expediency [598]. Experimental [626, 136, 137]. Explaining [661]. Explanation [648, 463, 128, 629, 533]. Explanations [553]. explanatory [614]. Explicative [20]. Explikation [291]. Extrapolation [90, 65].

Facts [612, 426]. Faktenkonstitution [206]. Faktentranszendente [187]. faktischer [388]. Fall [262]. Fallbeispiel [550]. Fallibilism [529, 461]. Fallibilismus [529, 461]. fallibilistischer [313]. Fallibility [621]. Fallstudie [257]. Fallstudien [148]. Falsifiability [345, 656]. Falsifiable [656, 602]. Falsification [175, 267]. Falsifikation [66, 210, 4]. Falsifikationismus [148]. Falsifikationsprinzip [152]. falsifizierbar [309]. Falsifizierbarkeit [185, 66]. Far [538]. Farbenlehre [402]. Fehlbarkeit [621]. Fehlschluß [451, 467]. Feigl [94]. Ferdinand [203]. Fertilization [535]. Festschriften [269]. Feyerabend [342, 355, 342, 489, 304]. Field [54, 522]. Fifth [360].

Figur [367]. Final [558]. Finalisierung [336]. Finalisierungsdebatte [301]. Finalization [440, 373, 578]. Finetti [513]. Finke [543]. Finland [8]. First [603]. FitzGerald [437]. Flecks [326]. fondements [521]. Forderung [371]. Foreword [1]. Formalismus [284]. Formation [169]. Formen [412, 32]. Formulierung [292]. Forschung [11, 49]. Forschungsbericht [451, 467]. Forschungspolitik [234]. Forschungsprogramme [150, 282]. Forschungsprogramms [230]. Forschungsprozessen [49]. forschungsstrategischer [376]. Forschungswissenschaft [37]. Fortpflanzung [659]. Fortschritt [549, 380]. Fortschritts [413, 151]. Foucault [439]. Foundation [559, 534, 577]. Foundational [360]. Foundations [510, 360, 639]. Four [360, 655]. Four-Color [655]. Fourth [464]. Frage [151]. France [51]. Francis [344]. Francisco [640]. Frank [216]. Franz [378]. Franzen [499]. Frederick [500]. Frege [648]. Freges [280, 648]. freie [342]. French [457, 327, 521, 51]. Freud [593]. Freundlieb [377]. Frey [53]. Fricke [377]. Friedrich [245, 246]. Frithjof [306]. Fritz [160]. Front [14, 183, 220, 254, 289, 324, 363, 400, 431, 455, 485, 517, 546, 591, 610, 643]. Frontiers [607]. Fruitful [493]. Function [329, 331]. Functional [316, 507]. Functionalization [373]. Fundament [245, 246]. fundamental [536]. Fundamentale [286]. Fundamentalismus [264]. Funktion [88, 415, 458, 133, 492]. Funktionalanalyse [460]. Funktions [507]. Funktions- [507]. Funktionsbestimmung [331]. Funktionsgeschichte [448]. Furniture [341].

Gadamer [40]. Gal [478]. Gal-Or [478]. game [596]. Gängelband [475]. ganz [633]. Ganze [54]. Ganzen [98]. Garcia [606]. Gavagai [599]. geben [335]. Gebhard [247]. Gebildes [250]. Gedächtnis [151]. Gedanken [331, 556]. Gedankenexperimente [118]. gegebenem [556]. gegen [209]. Gegensatz [209]. Gegenstand [122]. Gegenstandsbestimmung [284, 302]. gegenwärtige [24]. gegenwärtigen [536]. Gegenwartsphilosophie [52]. Gehalt [185, 550]. Geisteswissenschaften [359, 2, 378]. Geldsetzer [140, 269]. Geleitwort [1]. General [256, 566, 94, 119, 650, 550]. Generalized [651]. GenEthik. [659]. Genies [633]. Genius [633]. Geochronology [575]. Geochronometrie [115]. Geometrie [332, 469, 579, 654]. Geometrodynamik [115]. geometry [579, 654]. Georg [156]. Geowissenschaft [498]. Gerard [41, 481, 270]. Gerhard [53, 380, 480, 661]. Gerigk [250]. German [167, 2, 528, 633, 16, 13, 1, 56, 78, 95, 79, 76, 164, 201, 200, 236, 272, 307, 323, 343, 384, 416, 444, 468, 545, 561, 588, 623, 663, 17, 510, 281, 636, 245, 326, 63, 418, 233, 566, 567, 88, 648, 64, 190, 634, 560, 83, 222, 335, 578, 103, 109, 122, 291, 140, 185, 191, 230, 306, 413, 415, 147, 166, 257, 336, 419, 568, 503, 641, 582, 391, 494, 198, 271, 381, 482, 163, 402, 465, 519, 548, 352, 361, 392, 549, 362, 398, 429, 454, 483, 501, 608, 318]. German [451, 467, 148, 258, 520, 19, 11, 516, 531, 544, 589, 379, 469, 196, 263, 337, 309, 353, 153, 210, 130, 660, 292, 425, 452, 509, 54, 250, 377, 458, 543, 53, 171, 3, 31, 120, 165, 180, 442, 504, 635, 650, 26, 68, 158, 149, 141, 303, 340, 559, 603, 205, 286, 264, 251, 601, 150, 542, 662, 637, 276, 445, 459, 613, 113, 344, 172, 321, 199, 69, 571, 206, 476, 259, 386, 524, 477, 460, 192, 620, 473, 584, 260, 310, 487, 532, 595, 651, 151, 443, 435, 261, 300, 474, 285, 268, 404]. German [4, 27, 32, 52, 42, 241, 247, 114, 277, 346, 393, 397, 132, 115, 107, 106, 224, 579, 207, 614, 89, 208, 293, 339, 461, 621, 133, 246, 47, 55, 40, 72, 329, 212, 116, 98, 84, 436, 75, 142, 225, 284, 302, 342, 412, 514, 585, 5, 45, 90, 99, 117, 154, 152, 371, 134, 507, 414, 529, 628, 354, 226, 394, 209, 57, 50, 77, 278, 186, 615, 629, 533, 488, 616, 437, 33, 232, 294, 320, 387, 396, 617, 550, 574, 652, 653, 311, 203, 358, 85, 6, 70, 447, 551, 618, 34, 179]. German [410, 489, 101, 370, 388, 355, 552, 265, 587, 389, 227, 448, 213, 38, 238, 155, 86, 243, 37, 234, 313, 39, 124, 156, 269, 499, 282, 411, 586, 304, 450, 330, 580, 331, 65, 438, 248, 332, 515, 21, 10, 48, 66, 161, 195, 380, 481, 390, 470, 35, 74, 67, 640, 428, 244, 498, 266, 374, 607, 182, 219, 253, 287, 322, 218, 270, 659, 596, 619, 348, 22, 30, 317, 296, 356, 441, 420, 576, 525, 536, 658, 661, 280, 577, 110, 143, 349, 408, 526, 581, 654, 453, 597, 421]. German [475, 537, 262, 301, 333, 471, 553, 49, 297, 105, 556, 135, 160, 71, 170, 214, 539, 598, 283, 350, 375, 492, 121, 187, 357, 439, 554, 376, 472, 378, 557, 508, 422, 319, 188, 146, 369, 493, 24, 28, 177, 334, 513, 92]. Gernot [637]. Gesamttheorien [557]. Geschichte [2, 425, 303, 142, 203, 70, 439, 520]. Geschichtsphilosophie [195]. Geschichtswissenschaft [262]. Gesellschaft [425, 63, 554]. Gesellschaftstheorie [161, 508]. Gesetze [48]. Gesetzes [241]. Gesichtspunkt [379]. Gesichtspunkte [241]. Gespräche [54]. gewöhnlichen [633]. Gibt [557]. Gilbert [434]. Giltigkeit [99, 154, 155]. Gnoseological [366]. Gödel [233]. Gödels [214]. Gödelschen [170]. Goethes [402]. Gonseth [203]. Goodman [193]. Goodmans [192]. Göteborg [37]. Göttingen [396]. Gracia [639]. Grad [459, 613]. Gravitationstheorie [651]. Gravity [409, 651]. Grenze [98, 577]. Grenzen [452, 31, 620, 481, 607]. Grete [559]. Griaznov [25]. griechischen [135]. Größe [206]. Großer [333]. Grounding [503]. Grover [94]. Growth [73]. Grue [328]. Grundfragen [72]. Grundkurs [198, 271]. Grundlage [352, 378]. Grundlagen [510, 532, 474]. Grundlagenproblem [487]. Grundlagenstreit [488]. Grundlagenstreits [536, 311]. Grundlegung [19, 39, 577]. Grundprobleme [160]. Grundsatz [214]. Grundsatzfragen [445]. Grundzüge [435]. Guessing [424]. Guide [500]. Guidebook [500]. Guiding [223]. Gunnar [380, 481]. Günter [74]. Gurwitsch [130].

H [361]. Habermas [64, 339, 74]. Hacking [266]. Hackings [374]. hallmark [634]. Hamburg [396]. Handlung [39, 508, 319]. handlungstheoretischen [474, 508]. Handlungstheorie [286]. Hans [560, 465, 303, 585, 124, 332, 453, 177, 465, 382, 382]. Hans-Michael [303]. Hans-Ulrich [560, 585]. Happened [564]. Harald [377]. Harnatt [231]. Hartner [417, 428]. Hauptrichtungen [284, 302]. Heelan [174, 137]. Hegel [88, 412]. Hegel-Kongreß [412]. Heidegger [40, 238]. Heisenberg [202, 54, 173]. Heißen [524]. heißt [552]. Heitsch [320]. Helberger [251]. Helmut [286, 199, 55]. Helpful [211]. Hempel [372]. Henri [479]. Henry [559]. Henry-Hermann [559]. Herausforderung [270]. Herausgeber [79]. Herbert [94]. Hermann [270, 559, 142]. hermeneutic [331]. Hermeneutics [570, 603, 136, 228, 40, 450, 137]. Hermeneutik [149, 133, 47, 72, 448, 141, 132]. Hermeneutikkongreß [603]. Hermeneutische [3]. hermeneutischen [225, 331]. hermeneutischer [448]. heuristic [550]. Heuristik [458]. heuristischer [550]. Hierarchies [314]. Hillel [274]. Himmelsmechanik [310]. Hindsight [491]. Hinske [38]. Hintikka [360]. Hinweise [268]. Histoire [606, 327]. Historical [503, 360, 511, 12, 585, 514]. Historik [303]. historisch [225]. historisch-hermeneutischen [225]. Historische [206, 411, 514]. historischen [503, 585, 134]. historischer [206, 48]. History [520, 383, 329, 496, 407, 491, 228, 25, 327, 507, 653, 515, 548]. hoc [612]. Hoche [560]. Hochkeppel [270]. Hölling [106]. Hollis [252]. Holzkamp [355, 39]. Holzkamps [89]. Honorem [269]. Horst [250, 635, 635]. Hösle [633]. Hösles [633]. Hübner [270, 584]. Human [640, 100, 659, 127]. Humanism [347]. Humanistic [627]. Humberto [640]. Hume [168]. Hungarian [109]. Huygens [46, 60]. hypothèse [457, 613]. Hypothesen [459]. Hypotheses [308]. hypothesis [457, 613]. hypothetisch [3]. hypothetisch-deduktive [3].

Idea [585]. Idealisierte [553]. Idealism [633]. Idealismus [633]. Idealized [553]. Idee [425, 585]. Identität [404, 551, 524]. Identity [59, 569, 524, 551]. Ideologie [88]. Ideologiekritik [161]. If [424]. ihr [284, 302]. ihre [402]. ihrer [292, 276, 507, 179, 269, 35, 556]. ihres [571]. II [302, 200, 467, 60, 42, 77, 470]. III [140]. Illustrated [491]. im [636, 566, 222, 140, 582, 198, 271, 452, 54, 458, 442, 344, 443, 225, 488, 653, 332, 515, 244, 619, 170, 539, 350]. Imagination [446]. Immanent [423]. immanente [281]. Impact [500]. impliziert [191]. importance [525]. Imre [73, 151, 146]. Incommensurability [490, 87]. incompatible [557]. Inconsistency [406, 539]. independent [277]. Indeterminacy [628]. Indeterminiertheit [628]. Individualismus [311]. Individuality. [639]. induction [192]. Inductive [279, 519]. Induktion [5, 422]. Induktionsproblem [244, 135]. induktive [519, 135]. Inferences [279]. Infinitely [510]. infolge [619]. Information [15, 44, 81, 112, 145, 184, 221, 255, 290, 325, 364, 401, 432, 456, 486, 518, 547, 592, 611, 644, 619]. Informationsberges [619]. informationstechnischen [619]. Inkonsistenz [539]. Innate [279]. innovatives [49]. Inquiry [466]. Institut [37]. integrativen [49]. Intellectual [329]. Intelligence [555, 605]. intentionale [451, 467]. intentionalistische [451, 467]. Intentionalität [628, 615]. Intentionality [628, 534, 615]. Interest [605]. Intern [396]. International [360]. Internationale [452]. Internationalen [68, 158, 69, 476]. Interpretation [635, 40, 215, 576, 537]. Interpretationen [375]. interpretieren [537]. Intersystem [520]. Intersystemarer [520]. Invarianzgesetze [114]. Investigation [511]. Irrational [256]. Irrationale [453]. irreale [259]. Irrtum [474]. Isis [586]. istorii [25]. Italy [91]. Iterabilität [628]. Iterability [628]. Iterative [540]. IV [94]. Ixmann [599].

J [64, 543, 360, 397, 12, 640, 639]. Jahr [332]. Jahrhundert [140, 515]. Janich [153]. Janich/Kambartel/Mittelstraß [153]. Janichs [393]. Japan [158]. Japanese [36]. Jean [606, 385]. Jeffrey [500]. Joachim [300, 106]. John [197]. Jörg [498]. Jorge [639]. Jörn [303, 514]. Joseph [12]. Judgment [648]. judgments [653, 576]. Juhos [56, 45]. Juli [476]. jüngste [489]. Jürgen [250, 247, 107, 339, 74, 658]. Jurisprudenz [26]. Justification [528, 93, 541, 548, 461].

K. [211]. Kambartel [153, 246]. Kambartels [245]. Kamlah [218]. Kamlahs [392, 474]. Kamp [285]. kanadischer [163]. kann [335]. Kant [412, 637, 38, 168, 62, 598]. Kantian [653]. Kantische [653]. kantischen [243]. Kants [38, 598]. Karel [497]. Karl [567, 285, 305, 195, 567, 660, 315]. Kartenparadoxon [374]. Kausalanalyse [207]. Kausale [438]. kausalen [291]. Kausalgefüge [259]. Kausalität [566, 568]. Kausalitätstheorie [190]. Kennzeichen [634]. Keplerschen [310]. Kernlogik [116]. Kerridge [374]. Kibernetika [178]. Kirchberg [452]. Kirchberg/Wechsel [452]. Kirchgäßner [247]. Kisiel [138]. Klassische [418, 381]. Klaus [39, 658, 156]. Knowing [424]. Knowledge [503, 46, 60, 627, 424, 640, 169, 650, 619, 317, 73, 126]. Knowledge-Formation [169]. Kockelmans [12]. kognitiver [574]. Kollektivismus [311]. Kombinationstests [386]. Kommunikationsreflexion [542]. Komparative [34]. Konflikt [566]. Kongreß [68, 158, 69, 476, 412]. KongreßBerichte [95, 200]. Königs [370]. Können [191]. Konrad [357]. Konsenstheorie [64, 205]. Konsenstheorien [339]. konstituieren [116]. Konstitutionsproblem [285]. konstruieren [552]. Konstruktion [213]. konstruktive [294]. konstruktiven [579]. konstruktiver [188, 117]. Konstruktivismus [442]. Konstruktivistische [352]. konstruktivistischen [281]. Kontext [191]. Kontraktion [437]. Kontroverse [451, 467, 101]. Konventionalismus [115, 369]. Konventionalität [196]. Konvergenzen [52]. Konzepte [550]. Konzeption [122, 451, 467, 156]. Konzeptionen [567]. Kopenhagener [537]. Kopnin [11, 76]. Koppe [378]. Kordig [93]. Korrespondenz [205]. Korrespondenz- [205]. korrespondierenden [76]. Kraft [180, 404, 65]. Krafts [155, 165]. Kreativität [268]. Kriteriums [185]. Kritik [2, 465, 353, 509, 458, 206, 474, 397, 585, 489, 370, 280, 49, 281, 293, 410, 213]. kritisch [89]. kritisch-emanzipatorischen [89]. Kritische [258, 559, 150, 580, 88]. Kritischem [296, 101]. kritischen [63, 293, 74]. Kritischer [296, 230]. Kues [613, 613]. Kuhn [439]. Kuhnian [407]. Kuhnsche [394]. Kultur [332]. kultureller [586]. Kunstforschung [86].

Künstliche [471]. Kurt [584, 270, 659, 160]. Kybernetik [156]. Kyoto [158].

L [500]. Laboratory [426]. Lage [103]. Laird [235]. Lakatos [73, 151, 146, 465, 423, 356, 441, 493]. Language [403, 527, 551, 576]. language-analytical [576]. Lärm [301]. Larson [500]. Läßt [116]. Latour [426]. Lauener [479]. Law [374, 539, 128]. Laws [625, 277]. Leap [633]. Lehre [179]. Leibnizschen [310]. leistet [509]. Lenk [124]. Leonard [559, 396]. Lessing [585]. Letztbegründung [528, 529, 461]. Level [308]. Lexikon [361, 142]. L’homme [471]. Library [108]. Lichte [653, 244]. Life [136, 534, 426, 137]. Life-World [136, 534, 137]. light [653]. Likelihood [374]. likeness [615]. limit [577]. Limits [620]. Linguistic [530]. Linguistics [360]. Linguistics/Historical [360]. linguistische [476]. Link [462]. literarische [367]. literarischen [250]. Literary [466, 587, 530]. Literatur [471]. Literature [543]. literaturtheoretischen [284, 302]. Literaturwissenschaft [318, 284, 302, 377, 543, 587]. Lobkowicz [270]. Logic [360, 97, 108, 249, 61, 235, 556, 519, 601, 653, 521, 602, 597, 360]. Logical [522, 382]. logically [557]. Logik [191, 120, 601, 69, 476, 98, 447, 470, 556, 519, 11, 165, 321, 98, 653, 390, 597]. Logikbegründung [442]. Logikgeschichte [653]. Logique [521]. logisch [488, 557]. logisch-mathematischen [488]. logische [21]. logischen [271]. logischer [388]. Lorentz [495, 595, 437]. Lorentz-Kontraktion [437]. Lorentztransformationen [595]. Lorenz [357]. Lorenzen [27]. loss [536]. Löst [207]. Lösung [212, 580, 283]. Lothar [660]. Low [308]. Lübbe [270]. Ludwig [375, 487]. Ludwigs [350]. Ludwik [326]. Lutz [140, 269]. Lyssenkoismus [336]. Lyssenkos [411].

M [11, 341, 395]. Mach [567, 567]. machine [471]. Machines [638]. Macht [413]. Macroscopic [649]. Mai [425]. Mainzer [658]. Making [359]. Man [252, 512]. management [619]. Mann [333]. Mannheim [315]. Märchen [355]. Marcuse [74]. Maria [382]. Mario [123, 129]. Martin [252, 238]. Marx [88, 296, 427]. Marxian [338]. Marxismus [251]. Marxist [215]. marxistischen [251]. Maskerade [528]. Masquerade [528]. Masse [261]. Materie [469]. Mathematical [522]. Mathematics [510, 360, 655, 146]. Mathematik [203, 510, 320]. mathematisch [477]. mathematisch-biologischer [477]. mathematischen [488]. Matter [29, 14, 58, 80, 111, 144, 183, 237, 220, 273, 254, 288, 289, 324, 363, 399, 400, 430, 431, 455, 484, 502, 485, 517, 562, 546, 590, 591, 609, 624, 610, 642, 643, 323, 123]. Maturana [640]. Max [420]. Maxwell [94]. mean [552]. Meaning [564, 329, 628, 634, 277, 197]. Measurement. [497]. Mechanics [649, 536, 537]. Mechanik [471]. mechanization [659]. Mediation [556]. Medical [604]. Medicine [500]. Memoriam [239]. Menne [443]. Menschen [471, 342]. Menschliche [414]. menschlichen [122, 640]. menschlicher [659]. Mental [593, 634]. Mentalen [634]. Meßtheorie [435, 21]. Messung [397]. Meta [277]. Metagesetze [277]. Metaphilosophie [149]. Metaphysical [575]. metaphysics [521, 639]. Metaphysik [149]. métaphysique [521]. Metascience [41]. Method [405, 123]. Methode [251, 474, 354, 135, 88, 258, 3, 393]. Methodenlehre [53]. Methodenstreit [296]. Methodenzwang [304]. methodisch [352]. methodische [113]. Methodisches [27]. Methodological [338, 540, 175, 650]. Methodologie [150, 69, 476, 386, 443, 282]. Methodologien [548]. methodologies [548]. methodologische [4, 293]. methodologischen [2, 379, 650, 85]. Methodology [240, 612, 338, 360, 540, 622, 594, 215]. Methods [94, 530, 512]. metodologija [178]. Metrisierung [222]. Michael [94, 303, 239]. micro [536]. Miezekätzchen [355]. Mikrowelt [536]. Mikulinskij [25]. Mind [604]. Mind-Body [604]. Minnesota [94]. Mirror [450]. Misreadings [424]. Misuse [118]. Mitglieds [76]. Mitteilung [79]. Mittelstraß [153, 107]. modal [524]. Modalprobleme [524]. Model [20, 123, 229, 613]. Modellbegriff [348]. Modelle [49]. Modellplatonismus [247]. Models [646, 614]. Models. [427]. Modern [223, 615, 616, 617, 653]. moderne [52]. modernen [617, 653, 367]. moderner [615]. Modes [317]. Modification [365, 267]. Möglichkeit [33, 574]. Möglichkeiten [31]. Momente [222]. Monism [61]. Morscher [480]. Moskau [68]. mountain [619]. multiplicity [617]. Münch [161]. Münchausen [528, 596]. Münchhausen [528, 596]. Münchhausen-Trilemma [596]. Musgrave [73]. Myth [572, 584]. Mythos [584].

nach [371, 186]. Nachruf [130, 45, 30, 146]. Nachtrag [67, 296]. Nachträge [56, 663, 180]. Nagai [321]. Nagel [493]. Nagels [292]. Naming [551]. Natur [662]. Natural [569, 617, 539]. Naturbeschreibung [414]. Nature [625, 450, 662]. Naturgesetz [260, 421]. Naturgesetzlichkeit [539]. Naturwissenschaften [617]. naturwissenschaftlicher [459, 47]. nauki [178, 25]. Neemann [179]. Negative [535]. Negentropie [419]. Negro [52]. Nekrolog [209]. Nell [252]. nella [482]. Nelson [559, 396]. Nelsons [559]. Neo [252]. Neo-Classical [252]. ness [612]. network [619]. Neubegründung [542]. Neubestimmung [571]. neue [619, 75]. neuen [528, 353, 39]. neueren [499]. neuerer [448]. neues [246]. Neueste [340]. Neumaier [480]. Neuzeit [107]. neuzeitlichen [404]. neuzeitlicher [616]. nevertheless [557]. Newtonsche [260]. Nicholas [108, 162, 607]. nicht [537, 276]. Nicht-empirische [276]. Nichtexistenz [172]. nichtklassische [381]. nichts [301]. Niederlanden [92]. Nikolaus [613, 270, 613]. Nominalism [554]. Nominalismus [554]. nomologischer [226]. Non [510, 612, 394]. Non-ad-hoc-ness [612]. Non-standard [510]. Non-Statement [394]. Noosphere [631]. Norbert [38]. Nordenstam [125]. Norm [165]. normalen [492]. normativ [392]. normativ-analytischen [392]. Normative [627]. normativen [245, 246, 475]. Normen [154, 155]. North [511]. Norway [104]. Note [466, 298]. Novel [612]. nuova [482].

Objective [633, 126]. Objectivity [315, 527, 216]. objektiven [633]. Objektivität [167]. Ocerki [25]. ohne [566, 425]. Ökologie [550]. Ökonomie [245, 251, 247, 246, 246]. Oktober [396]. O’Malley [197]. One [360]. ontologic [537]. Ontological [61]. Ontologie [262, 278]. ontologisch [537]. Operation [96]. Operative [241]. Ordinary [633]. Organic [604, 598]. organischer [598]. Origin [457]. Origine [457]. Orman [479]. Ortsbestimmung [452, 134, 507]. Österreich [24]. Österreichischer [28]. Oswald [662]. Otto [480]. Our [279]. Outline [446]. overall [557]. overcome [616]. Overcoming [559]. Own [655].

P [76, 190, 11, 601, 393, 305]. P. [601]. Pädagogik [571]. pädagogische [179]. pädagogischen [370]. Palmer [40]. pankritischen [116]. Paolo [482]. Paradigm [573, 555]. Paradigmawechsel [83]. Paradigms [82]. Paradox [326, 193]. Paradoxes [602]. Paradoxon [147]. paribus [345]. Parity [523]. Part [54]. Parteilichkeit [167]. Partial [82]. Particle [223, 523]. Particulars [59]. Parts [360]. path [601]. Patrick [137]. Paul [27, 342, 489, 500, 304, 275]. pedagogy [571]. Peirce [521]. Permanenz [371]. Persistent [424]. Personal [496]. Perspectivity [534]. Perspektive [477]. Perspektiven [70, 344]. Peter [415, 543, 128]. Peukert [286]. Pfarr [300]. Phänomenologie [615]. phänomenologischer [130]. Phase [423]. Phenomena [649]. Phenomenology [505, 615]. Philological [549]. Philologie [352]. Philologischer [549]. Philos [396]. Philosophen [536, 349]. Philosopher [146]. philosophers [109, 163, 536]. Philosophical [252, 360, 530, 619]. Philosophie [140, 452, 340, 542, 69, 42, 72, 412, 618, 10, 619, 408, 471, 105, 51, 327, 560, 83, 402, 53, 31, 149, 386, 652, 243, 658, 480, 297, 135, 542]. Philosophiegeschichte [140]. Philosophiehistorie [507]. Philosophieren [637]. Philosophies [139]. philosophische [233, 349]. philosophischen [321, 619]. philosophischer [72, 269]. Philosophy [91, 18, 157, 194, 604, 94, 360, 542, 181, 176, 8, 159, 249, 534, 36, 104, 305, 450, 128, 217, 23, 9, 578, 327, 582, 571, 507, 615, 629, 652, 618, 619, 577, 51, 560, 637, 658, 480, 542, 108, 478, 383, 12, 129]. Physical [464, 277]. Physicist [18]. Physics [96, 433, 522, 223, 523, 94, 572, 478, 658, 462, 551, 577, 537, 129]. Physik [402, 142, 658, 166, 196, 52, 447, 551, 475]. physikalische [414]. physikalischen [387]. physikalischer [277]. physiology [652]. Piaget [606, 385]. Plädoyer [413, 186, 537]. Planetary [491]. Planung [124]. plea [537]. Pluralismus [371, 199]. Poetic [530]. point [533, 574]. Poland [249]. Polanyi [239]. Polen [71]. Policy [359]. Polikarov [379, 540]. Politikwissenschaft [441]. politische [230, 391, 270, 356, 245]. Politischen [494, 251]. Popowitsch [11]. Popper [567, 305, 375, 365, 567, 185, 660, 338, 523, 126, 656, 358, 424, 602].

Popper-Diskussion [358]. Popper-Kriteriums [185]. Popperianischen [389]. Poppers [350, 208]. Porter [500]. Positionen [587]. positioning [507]. Positions [587]. Positive [34]. positiven [133]. Positivismus [284, 32, 369]. Positivismusstreit [67]. Positivist [564]. Posner [530]. possibilities [507]. possibility [574]. Possible [315]. Practical [528, 542]. pragmatic [636, 614, 461, 629]. Pragmatics [542, 506, 331]. Pragmatik [542, 331]. pragmatische [636, 422, 629]. pragmatischen [614]. pragmatism [521]. pragmatisme [521]. praktischen [542]. praktischer [528]. Praxis [132]. Predictability [566]. Primat [98]. Principle [366]. Principles [223]. Prinzipielle [234]. priori [576]. Priority [573]. Privatrechts [26]. Probability [513]. Problem [565, 647, 264, 259, 115, 5, 313, 580, 48, 35, 462, 539, 598, 330, 604]. Problematik [171, 276]. Probleme [42, 488, 70, 659, 379, 206, 278, 179, 10]. Problems [647, 360, 659, 583, 524, 497]. Problemstellungen [340, 234]. Procedures [403, 522]. Proceedings [360]. Professor [578, 138, 657]. Prognose [124]. Prognosen [472]. program [465]. Programm [465]. Programme [575]. Programmes [612]. Programs [523]. Progress [490, 424, 549]. Progressive [523]. Projects [240]. Projektionsoperatoren [224]. Proof [655]. propensities [350]. propensity [208]. propensity-Theorie [208]. Propositions [528]. Protophysik [281, 300, 474, 393, 410, 227, 475, 397]. Prozeß [222]. Prüfbarkeit [504, 635, 581]. Prüfung [148]. Psychogenèse [606]. Psychologie [89, 574, 521, 6]. psychologism [597]. Psychologismus [597]. Psychology [462, 521, 574, 94]. Publications [109, 163]. Pursuit [405]. Puzzles [97].

Quantenmechanik [537, 283, 536, 375]. Quantenphysik [537, 577]. Quantentheorie [487]. Quantification [61]. Quantifizierung [376, 222]. quantitativen [222]. Quantum [409, 522, 649, 536, 577, 537]. Quest [347]. Quine [479, 628, 657, 628].

R [93, 660, 141, 360, 126, 500, 450, 640, 25]. rôles [457]. Radder [465]. Radermacher [361]. Radical [173, 529]. radikaler [529]. Radner [94]. Radnitzky [380, 481, 270, 41]. ragione [482]. Rapp [298]. Rational [256, 466, 329, 548, 252, 530]. rationale [548, 458]. Rationalismus [230, 293, 116, 101, 296]. Rationalität [380, 150]. Rationality [82]. ratlos [536]. Rätsel [192]. Raum [397, 658]. Raum-Zeit [658]. Rautenberg [381]. razvitija [25]. Re [315]. Re-Evaluation [315]. Realisierbarkeitspostulates [191]. Realisierungsbedingungen [598]. Realismus [342, 67, 106, 101]. realistic [537]. realistic-ontologic [537]. realistisch [537]. realistisch-ontologisch [537]. realistischen [64]. Realität [437]. Reality [495, 464]. realization [598]. reason [542, 576, 514, 585]. reason-critical [576]. Reasoning [529, 539]. received [545, 561, 623]. Rechenverfahren [390, 470]. Rechtfertigung [422]. Rechtstheorie [121]. Reconstruction [329, 548]. Reconstructive [542]. redefine [571]. Reduction [572]. Reductionism [462]. reduktionistischer [292]. Refinements [242]. Reflection [542, 505, 534]. Reflections [559, 508]. reflective [507, 529]. Reflexionsmöglichkeiten [507]. Reflexive [529, 214]. Refounding [542]. Regeln [113, 293]. Regulativ [615]. regulator [615]. Rehabilitation [510]. Rehabilitierung [510]. Reichenbach [382]. Reichweite [452]. Reihe [27]. Reihen [75]. Rejection [267]. Rekonstruktion [548, 436, 311]. Rekonstruktive [542]. related [539]. Relationism [315]. Relations [449]. Relationship [405, 571, 615, 618, 586]. relativ [408]. Relativismus [151]. relativistischen [21]. Relativität [106]. Relativitätsprinzip [300, 227]. Relativitätstheorie [566, 532]. Relativity [365, 566, 532, 406]. Relevanz [120]. Reliability [279]. Remarks [131, 522, 601, 600, 580]. Renormalization [522]. Repeated [308]. Replica [529]. Replik [529]. Reply [231, 174, 657, 211]. Representations [427]. Reproducibility [550]. reproduction [659]. Reproduzierbarkeit [550]. Rescher [108, 162, 607]. Research [240, 612, 403, 46, 60, 131, 223, 523, 36, 312, 575, 490, 359, 491, 189, 334]. response [465]. Responsibility [565]. result [619]. Resultate [42, 10, 318]. results [550]. Rettung [448]. Review [558, 73, 93, 252, 54, 604, 94, 360, 383, 622, 108, 178, 497, 126, 606, 12, 41, 479, 478, 305, 162, 500, 450, 530, 125, 127, 128, 197, 217, 235, 216, 605, 638, 639, 123, 129, 341, 395, 427, 512, 583, 25, 382, 426, 509, 636, 560, 660, 543, 559, 542, 662, 637, 584, 514, 587, 586, 515, 640, 607, 659, 658, 661, 513]. Review-Discussion [450]. Reviews [585]. Revisited [573]. Revolte [89]. Revolutionen [147, 263]. Revolutions [409, 242]. Rezension [636, 560, 140, 306, 413, 415, 198, 271, 381, 482, 361, 11, 379, 196, 660, 54, 250, 543, 53, 26, 141, 303, 559, 286, 251, 542, 662, 637, 321, 199, 477, 584, 443, 285, 27, 52, 42, 397, 107, 106, 55, 40, 72, 75, 142, 514, 414, 320, 396, 179, 587, 38, 39, 124, 269, 499, 586, 304, 515, 10, 161, 195, 380, 481, 74, 640, 498, 607, 218, 270, 659, 658, 480, 661, 453, 105, 160, 378, 319, 177, 513]. Rezensionen [377, 342, 585]. Richard [40, 161, 604]. Richtigkeit [339]. riddle [192]. Rise [119]. Robert [382]. Rohrmoser [74]. Roland [530]. Rolando [606]. roles [457]. Romania [194]. Ronal [500]. Roots [405, 640]. Roper [500]. Rorty [450]. Rossi [482]. Rossini [500]. Rüdiger [319]. Rudjer [519]. Rudolf [285, 30]. Rüsen [303, 514]. Russells [280].

S [521, 38, 500, 25, 382]. Sache [292]. Salmon [382]. Salzburg [476]. Sätze [528]. Savigny [377, 198, 271]. Schäfer [660]. Schaffens [6]. Schein [192]. Schein-Rätsel [192]. Schema [365, 523]. Schilpp [305]. Schleiermacher [40]. Schließen [271]. Schmidt [543]. Schools [41]. Schriften [56, 588, 663, 180]. Schurz [661]. Schweiz [50, 203]. schweizerischer [57, 77]. Schwemmer [662]. Schwierigkeiten [121]. Science [91, 565, 647, 157, 434, 520, 194, 338, 94, 360, 594, 181, 176, 511, 136, 8, 159, 506, 249, 137, 139, 496, 12, 119, 572, 533, 295, 405, 36, 104, 242, 500, 102, 527, 538, 632, 23, 7, 25, 9, 229, 578, 109, 163, 548, 571, 615, 629, 618, 515, 525, 51, 567, 559, 540, 571, 607, 146, 661, 360, 622, 128, 216]. Sciences [204, 569, 360, 542, 315, 100, 51, 650, 617, 331, 619, 606]. Sciences/Basic [360]. Scientific [409, 93, 316, 612, 403, 625, 97, 405, 541, 490, 347, 602, 480, 427, 189, 334, 426, 567, 613, 533, 616, 550, 552]. scientific-theoretical [533]. Scientification [594]. Scienza [482]. Scope [620, 577]. Search [622]. Searle [628, 628]. Second [249]. security [621]. Seelig [414]. Sehweise [49]. Seiffert [55]. Sein [524]. seine [580, 598]. seiner [191]. Selbststudium [198, 271]. Selected [382]. Selective [181, 139]. Self [505]. Self-Reflection [505]. selon [521]. Semantics [570]. Semantik [460]. Semiconductor [407]. Seminar [303]. semiotische [185]. sensory [652]. September [603, 69]. Serien [95, 200]. Shen [603]. Shiro [321]. sich [116]. Sicherheit [621]. Sicht [340, 367, 471, 574]. side [331]. Siegfried [543]. significance [650]. Simulation [533]. Sinn [88, 280]. Sinnbedingungen [35]. Sinne [350]. Sinnesphysiologie [652]. Sittlichkeit [396]. Situation [24]. skeptischer [313]. Sketch [240, 534]. Skizze [230, 304]. Skizzen [124]. Small [510]. Social [256, 204, 569, 338, 622, 329, 216, 315, 426]. società [482]. Society [347, 554, 508, 235]. Sociological [317]. Sociology [405, 197]. Sociotectonics [631]. sog [192]. sogenannten [2, 397]. ’Sollen [191]. Solution [604, 580, 598]. Solving [647]. Some [131, 215, 602, 556]. Sorts [450]. sowjetischer [340]. Sozialforschung [376]. sozialphilosophischen [230]. sozialwissenschaftlichen [311]. sozialwissenschaftlicher [445]. Soziologie [67, 420]. Space [658]. Space-Time [658]. Speak [424]. Special [365, 360]. Speculation [61]. speculative [652]. spekulative [652]. Spinner [199]. sprachanalytische [408]. sprachanalytischen [576]. sprachanalytischer [349]. Sprache [377, 447, 551, 218, 349, 319, 477, 378]. Spracherlernung [477]. Sprachkritische [473]. sprachphilosophischen [378]. Sprachtheorie [285]. Sprachwissenschaftstheorie [285]. Spreer [306]. Sprung [633]. Squared [351]. Stand [536]. standard [510]. Standpunkt [116]. Stanley [500]. State [159]. Statement [394]. Statements [656, 602]. States [593]. statistische [166, 166]. statistischen [190]. status [536]. Stegmüller [351, 263, 42, 10]. Stegmüllerschen [266]. Steinvorth [494]. Stellung [332, 21]. Stephen [567, 94, 622, 127, 567]. Steve [426]. Stolpersteine [601]. störenden [172]. Story [407]. Stoßmechanik [418]. Strategie [292]. Strategien [318]. Ströker [515]. Strube [560]. Structural [569]. Structuralist [580, 635]. Structure [409, 464, 508]. Struktur [415, 147, 472, 508]. Strukturalismus [207, 302]. strukturalistische [580, 86]. strukturalistischen [635]. Struve [635]. Stüben [415]. Stück [203]. Studie [262]. Studien [140]. Studies [223, 94, 25, 525, 543, 587]. Study [62, 512, 550, 618]. Stumbling [601]. Stupidities [450]. Stuttgart [412]. Stützungsbegriffs [266]. Success [491]. Successor [449]. Suche [186]. Suggestions [507]. Superconductivity [649]. Superfluidity [649]. Superlative [34]. Suppes [190]. Supplements [663]. Survey [139, 9]. Symmetrie [210, 4, 617]. Symmetriebrechung [617]. Symmetry [175, 617, 617]. Sympathy [367]. Symposium [425, 452, 396]. Synthetic [125, 576]. synthetisch [576]. System [225, 420, 49, 462, 582, 619, 598]. system-theoretical [598]. Systematischer [451, 467]. Systembegriff [26]. Systemdenken [26]. Systems [565, 550, 395]. Systems-Theoretic [565]. systemtheoretische [598]. Systemtheorie [550]. Systemvergleich [582]. Szientismus [247].

T [500]. table [653]. Talks [54]. task [619]. Taxonomy [626]. technical [650]. Technik [413, 16, 650]. Technikphilosophie [582]. technischen [650]. Technisierung [659]. Technology [647, 500, 583, 619, 582, 650]. Teil [451, 467, 54, 284, 302]. Teleologie [419, 292, 425]. teleologische [438]. teleonomische [438]. Telos [425]. Temporal [108]. Tendenzen [340]. teorii [25]. terms [277]. testability [504, 635, 581]. Testing [96, 308, 20]. Tests [308, 626]. Teststrenge [389]. Tetradic [365, 523]. Texte [585]. Texts [585]. Textwissenschaften [451, 467]. Their [646, 571, 507]. Themen [619]. Theodor [74]. Theologie [286]. Theorem [655]. Theoretic [565, 173]. Theoretical [626, 580, 462, 533, 598, 508, 504]. theoretische [387, 504]. theoretischem [371]. theoretischen [580]. theoretischer [261]. Theorie [245, 63, 230, 250, 303, 27, 132, 208, 387, 370, 552, 243, 74, 296, 327, 277, 2, 391, 337, 595, 42, 394, 498, 356]. Theoriebildung [445]. Theorien [335, 309, 635, 504, 581]. Theorienbewertung [313]. Theorienwandel [515]. Theories [646, 316, 625, 522, 94, 497, 119, 440, 449, 504, 635, 515, 581, 557]. Theory [564, 256, 365, 520, 46, 60, 131, 360, 627, 651, 61, 40, 329, 446, 534, 20, 423, 162, 406, 102, 25, 277, 550, 618, 552, 508, 327, 595, 513]. Theory/Foundational [360]. There [359, 557]. These [566]. Thesen [619, 63]. theses [619]. Thesis [566, 59, 174]. Thoughts [331, 556]. Three [360]. Tiller [500]. Time [658]. Tode [165]. Todestages [396]. Tokyo [158]. Tool [61]. Topics [619]. Topitsch [177]. Tore [125]. Toulmin [567, 127]. Tradition [657]. Traditions [490]. Tragweite [620, 577]. Transcendental [528, 461, 577]. transformations [595]. transzendentale [101]. transzendentaler [528, 415]. Transzendentalphilosophie [38, 577]. transzendentalpragmatische [461]. Tree [640]. Trends [176, 36]. trilemma [596, 596]. truth [584, 615, 162]. Truth-likeness [615]. turn [636, 629]. Turner [622]. Two [360, 449, 630, 334]. Types [449, 334]. Typology [189].

Überblick [443]. Überlegung [576]. Überlegungen [122, 292, 425, 452, 404, 508]. übertragbar [469]. Überwindung [559, 247, 616, 85]. Übungen [198, 271]. UdSSR [76, 70]. Uhren [147]. Uhren- [147]. Ulrich [560, 585]. Ultimate [528, 529]. Umkreis [54]. Umrisse [303]. unabhängige [277]. unaufhebbare [98]. Unbeweisbarkeitsaussagen [214]. Understanding [403, 127, 7, 661, 526, 634, 427]. Unendlichkleinen [510]. Unentscheidbarkeit [233]. ungarischer [109]. Ungarn [103]. uniqueness [579]. Unity [628, 572, 617]. Universal [656, 632]. Universalität [330].

Universität [37]. Unterscheidung [222, 388, 280]. Untersuchung [276]. Untersuchungen [251, 285]. unveiled [586]. unvereinbare [557]. Unvollständigkeitsproblem [214]. Unvollständigkeitsproblems [170]. Urquhart [108]. Ursula [179]. Urteile [576]. Urteils [648]. Urteilstafel [653]. USA [139]. Use [118, 127]. using [571].

V [178, 25, 52]. Vagueness [372]. Vaihinger [332]. validation [613]. Varela [640]. Verabschiedung [448]. Verallgemeinerte [651]. verbesserter [226]. Verbund [619]. verfahrenden [352]. Vergleich [355, 375, 439]. Vergleichende [357]. Verhalten [122]. Verhaltensforschung [357]. Verhältnis [258, 310, 615, 618, 586, 475]. Verhältnisses [571]. Verification [175]. Verifikation [210, 4]. vermeintliche [153, 212]. Vermittlung [556]. Vernuft [396]. Vernunft [542, 370, 514, 585, 218, 319]. vernunftkritischen [576]. Veröffentlichungen [109, 163, 57, 77, 28]. Version [489]. Verstehen [2, 661, 204, 209]. verstehenden [420]. Verstehensbegriffs [526]. Versuch [528, 222, 292, 571, 116, 311, 39]. versus [461, 529, 450, 188]. Verwirrungen [633]. Verwissenschaftlichung [318]. Verzeichnis [76]. Victor [165, 180, 155, 65]. Viel [301]. Vielheit [617]. View [328, 496, 394, 635, 533, 574]. viewed [533]. Viewpoint [462]. Views [540]. Violation [523]. vis [359]. vis-à-vis [359]. Vision [469]. Vol [94, 108, 127]. Volker [477]. vollständig [537]. Volume [15, 44, 81, 112, 145, 184, 221, 255, 290, 325, 364, 401, 432, 456, 486, 518, 547, 592, 611, 644]. vom [122, 379, 425, 452, 68, 158, 69, 476, 116, 396, 37, 332]. Voraussetzung [404]. Voraussetzungen [64, 481]. vorgeschlagenen [375]. Vorhersagbarkeit [566]. Vorschlag [89]. Vorschläge [507]. Vorstellung [414]. Vorurteile [488]. vs [572, 550].

W [76, 11, 500, 74, 105, 427]. Wahn [488]. wahr [499]. wahre [332]. Wahrheit [584, 339, 499, 171, 205, 33, 447, 388, 244]. Wahrheitsähnlichkeit [615]. Wahrheitsbegriff [499]. Wahrheitskriterien [186]. Wahrheitstheorien [499]. Wahrscheinlichkeit [166, 353, 208]. wahrscheinlichkeitstheoretischen [283]. Wahrscheinlichkeitstheorie [294, 513]. Walter [52]. Wandels [567, 619]. Wandschneiders [213]. War [249, 333]. Wartofsky [427]. Was [355, 509, 552, 421]. Weber [296]. Webers [420]. Weberschen [241]. Wechsel [452]. Weg [601, 38]. Wege [436]. Wellbery [587]. Welt [557]. Weltanschauung [489, 320]. Wende [636, 476, 412, 629]. Werner [560, 196, 54, 202]. Wertfreiheit [296]. Werturteile [155]. Werturteilsfrage [16]. Werturteilsstreit [177]. Wesley [382]. Whatever [564]. Whole [54]. Wider [304]. Widerlegung [292]. Widersprüchlichkeit [192]. ‘Widerstreit’ [636]. Wie [335, 346]. Wiehl [141]. Wilhelm [26, 585, 218]. Willard [479]. Wille [414]. Willy [428, 270, 417]. Winfried [499]. Winokur [94]. Wirklichkeit [414, 203, 447]. Wirkungspotential [448]. Wirtschaftsplanung [306]. Wirtschaftswissenschaften [492]. Wissens [503]. Wissenschaft [76, 122, 336, 494, 53, 330, 480, 71, 492, 17, 122, 413, 559, 571, 404, 167, 39, 380, 481, 607, 270, 22, 661, 160, 71]. Wissenschaft/Essays [480]. Wissenschaften [425, 542, 69, 151, 296, 650, 133, 225, 342, 134, 331]. Wissenschaftler [453, 394]. Wissenschaftlers [367]. Wissenschaftliche [10, 263, 443, 552, 155, 268]. wissenschaftlichen [76, 567, 198, 11, 379, 31, 150, 613, 6]. wissenschaftlicher [147, 550, 265, 282, 472]. Wissenschaftsauffassungen [616]. Wissenschaftsbegriff [188]. Wissenschaftsbegriffes [19]. Wissenschaftsentwicklung [567]. Wissenschaftsethik [344, 618]. Wissenschaftsforschung [326, 618, 525]. Wissenschaftsgeschichte [425, 68, 158, 84, 515, 376, 548, 258]. Wissenschaftsgeschichtliche [83]. wissenschaftsgeschichtliches [330]. Wissenschaftskonzeption [355]. Wissenschaftskritik [153]. Wissenschaftslehre [367, 6, 39]. Wissenschaftsphilosophie [203, 578, 415]. wissenschaftspraktische [268]. Wissenschaftspsychologie [122]. Wissenschaftssystem [332]. wissenschaftstheorethischen [276]. Wissenschaftstheoretiker [57, 77, 109, 163, 130, 28]. wissenschaftstheoretisch [533]. Wissenschaftstheoretische [83, 251, 285, 342, 357]. wissenschaftstheoretischen [507, 411]. wissenschaftstheoretischer [222]. Wissenschaftstheoretisches [361]. Wissenschaftstheorie [95, 200, 326, 83, 103, 140, 306, 258, 520, 153, 377, 340, 571, 476, 8, 42, 346, 412, 618, 37, 10, 441, 23, 51, 24, 391, 494, 392, 55, 84, 152, 50, 615, 629, 85, 389, 37, 313, 124, 471, 92]. Wissenschaftstheorie. [286]. Wissenschaftswissenschaft [70]. Wissensordnung [619]. Wissenswissenschaften [619]. without [566]. Wittgenstein [452]. Wittgenstein-Symposium [452]. Wittgensteinian [574]. Wittgensteinscher [574]. Wolf [498]. Wolfgang [413, 381, 42, 414, 354, 10]. Wolfram [320]. Woolgar [426]. Work [131, 424]. World [136, 249, 534, 395, 536, 557, 341, 137]. Writings [382, 588, 663]. Würdigung [509]. Wurzeln [640].

XIX [425].

Yehoshua [274]. Yirmiahu [383]. York [477]. Yovel [383].

Z [142]. Zahn [601]. Zecha [480]. Zeit [281, 300, 397, 410, 227, 393, 658]. Zeitmessung [232]. Zeitmetrik [114]. Zhen [603]. Zimmermann [498]. Zitatenanalyse [525]. Zivilisationskritik [542]. Zufalls [559]. Zukunft [90, 65]. Zum [165, 115, 586, 35, 536, 245, 198, 271, 258, 337, 425, 635, 69, 151, 285, 246, 412, 615, 499, 67, 296, 172, 224, 152, 48]. Zur [2, 191, 230, 650, 149, 276, 617, 85, 101, 576, 378, 56, 95, 200, 663, 326, 63, 190, 122, 140, 415, 451, 467, 148, 180, 442, 251, 601, 150, 542, 206, 473, 310, 404, 436, 585, 628, 358, 38, 156, 218, 619, 577, 475, 262, 49, 556, 214, 376, 281, 306, 19, 353, 210, 377, 171, 635, 205, 386, 460, 151, 247, 397, 614, 371, 134, 616, 574, 410, 388, 265, 266, 170].

Zusammenhang [224, 170, 539]. Zustandekommen [88]. Zweckmäßigkeit [425, 598]. Zwei [148, 436]. zweite [260]. zweiten [393]. zwischen [210, 264, 224, 101, 355, 296, 280].

References

Anonymous:1970:GGF

[1] Anonymous. Geleitwort. (German) [foreword]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(1): 1, March 1970. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170165.

Albert:1970:TVG

[2] Hans Albert. Theorie, Verstehen und Geschichte: Zur Kritik des methodologischen Autonomieanspruchs in den sogenannten Geisteswissenschaften. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(1):3–23, March 1970. CO-DEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170166.

Frey:1970:HHD

[3] Gerhard Frey. Hermeneutische und hypothetisch-deduktive Methode. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(1):24–40, March 1970. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170167.

Juhos:1970:MSV

[4] Béla Juhos. Die methodologische Symmetrie von Verifikation und Falsifikation. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(1):41–70, March 1970. CO-DEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170168.

Kraft:1970:PIG

[5] Victor Kraft. Das Problem der Induktion. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(1):71–82, March 1970. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170169.

Mikulinskij:1970:PWS

[6] S. R. Mikulinskij and M. G. Jarosevskij. Psychologie des wissenschaftlichen Schaffens und Wissenschaftslehre. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(1):83–103, March 1970. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170170.

Yildirim:1970:TUS

[7] Cemal Yildirim. Towards an understanding of science. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(1):104–118, March 1970. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170171.

Hintikka:1970:PSW

[8] Jaakko Hintikka. Philosophy of science (wissenschaftstheorie) in Finland. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(1):119–132, March 1970. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170172.

Zeman:1970:PSE

[9] Vladimir Zeman. The philosophy of science in Eastern Europe a concise survey. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(1):133–141, March 1970. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170173.

Schleichert:1970:RPR

[10] Hubert Schleichert. Rezension: Probleme und Resultate der Wissenschaftstheorie und Analytischen Philosophie, Band I: Wissenschaftliche Erklärung und Begründung von Wolfgang Stegmüller. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(1):142–150, March 1970. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170174.

Diemer:1970:RLW

[11] Alwin Diemer. Rezension: Logik der wissenschaftlichen Forschung von P. W. Kopnin, M. W. Popowitsch. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(1): 150–152, March 1970. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170175.

Konig:1970:BRP

[12] Gert König. Book review: Philosophy of Science: The Historical Background by Joseph J. Kockelmans. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(1):152–154, March 1970. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170176.

Anonymous:1970:BG

[13] Anonymous. [Bibliographie]. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(1):155– 160, March 1970. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170177.

Anonymous:1970:FM

[14] Anonymous. Front matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(1):??, March 1970. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170164.

Anonymous:1970:VI

[15] Anonymous. Volume information. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(1):??, March 1970. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170163.

Albrecht:1970:WTG

[16] Ulrich Albrecht. Die Werturteilsfrage in der Technik. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(2):161–172, September 1970. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170179.

Apel:1970:WEG

[17] Karl-Otto Apel. Wissenschaft als Emanzipation?. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(2):173–195, September 1970. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170180.

Bunge:1970:PP

[18] Mario Bunge. The physicist and philosophy. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(2):

196–208, September 1970. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170181.

Diemer:1970:GAW

[19] Alwin Diemer. Zur Grundlegung eines allgemeinen Wissenschaftsbegriffes. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(2):209–227, September 1970. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170182.

Martin:1970:EMT

[20] Michael Martin. An explicative model of theory testing. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(2):228–242, September 1970. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170183.

Schleichert:1970:LSR

[21] Hubert Schleichert. Über die logische Stellung der relativistischen Meßtheorie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(2):243–251, September 1970. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170184.

Stegmuller:1970:WEG

[22] Wolfgang Stegmüller. Wissenschaft und Erklärung. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(2):252–263, September 1970. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170185.

Witt-Hansen:1970:PSW

[23] Johs. Witt-Hansen. Philosophy of science (wissenschaftstheorie) in Denmark. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(2):264–283, September 1970. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170186.

Zecha:1970:GSW

[24] Gerhard Zecha. Die gegenwärtige Situation der Wissenschaftstheorie in Österreich. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(2):284–292, September 1970.

CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170187.

Zeman:1970:BRO

[25] Vladimir Zeman. Book review: Ocerki istorii i teorii razvitija nauki (Studies on the History and Theory of the Development of Science) by V. S. Bibler, B. S. Griaznov, S. R. Mikulinskij. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(2): 293–295, September 1970. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170188.

Geldsetzer:1970:RSS

[26] Lutz Geldsetzer. Rezension: Systemdenken und Systembegriff in der Jurisprudenz, entwickelt am Beispiel des deutschen Privatrechts von Claus-Wilhelm Canaris. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(2):296–303, September 1970. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170189.

Juhos:1970:RMD

[27] Béla Juhos. Rezension: Methodisches Denken (Reihe Theorie 2) von Paul Lorenzen. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(2):304–310, September 1970. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170190.

Zecha:1970:VOW

[28] Gerhard Zecha. Veröffentlichungen Österreichischer Wissenschaftstheoretiker. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(2):311–321, September 1970. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170191.

Anonymous:1970:BM

[29] Anonymous. Back matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 1(2):??, September 1970. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170192.

Stegmuller:1971:NRC

[30] Wolfgang Stegmüller. Nachruf auf Rudolf Carnap. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(1):1–13, March 1971. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170195.

Frey:1971:MGW

[31] Gerhard Frey. Möglichkeiten und Grenzen einer wissenschaftlichen Philosophie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(1):14–26, March 1971. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170196.

Juhos:1971:FPG

[32] Béla Juhos. Formen des Positivismus. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(1):27–62, March 1971. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170197.

Lorenzen:1971:BMD

[33] Hans-Peter Lorenzen. Bemerkung über eine Möglichkeit der Definierbarkeit von Wahrheit. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(1):63– 65, March 1971. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170198.

Morscher:1971:PKS

[34] Edgar Morscher. Über Positive, Komparative und Superlative. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(1):66–88, March 1971. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170199.

Schnadelbach:1971:DEP

[35] Herbert Schnädelbach. Dispositionsbegriffe der Erkenntnistheorie: Zum Problem ihrer Sinnbedingungen. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(1): 89–100, March 1971. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170200.

Nagai:1971:RTJ

[36] Hiroshi Nagai. Recent trends in Japanese research on the philosophy of science. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(1):101–114, March 1971. CODEN ZAWTA2.

ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170201.

Radnitzky:1971:WFB

[37] Gerard Radnitzky, Håkan Törnebohm, and Göran Wallén. Wissenschaftstheorie als Forschungswissenschaft: Bericht vom Institut für Wissenschaftstheorie an der Universität Göteborg. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(1):115–119, March 1971. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170202.

Pivcevic:1971:RKW

[38] Edo Pivcević. Rezension: Kants Weg zur Transzendentalphilosophie: Der dreißigjährige Kant, 172 S. von Norbert Hinske. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(1):120–121, March 1971. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170203.

Rapp:1971:RWH

[39] Friedrich Rapp. Rezension: Wissenschaft als Handlung. Versuch einer neuen Grundlegung der Wissenschaftslehre von Klaus Holzkamp. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(1):122–129, March 1971. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170204.

Kisiel:1971:RHI

[40] Theodore Kisiel. Rezension: Hermeneutics: Interpretation Theory in Schleiermacher, Dilthey, Heidegger and Gadamer von Richard E. Palmer. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(1):130–132, March 1971. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170205.

Konig:1971:BRC

[41] Gert König. Book review: Contemporary Schools of Metascience by Gerard Radnitzky. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(1):133–137, March 1971. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170206.

Juhos:1971:RTE

[42] Béla Juhos. Rezension: Theorie und Erfahrung. Probleme und Resultate der Wissenschaftstheorie und Analytischen Philosophie, Band II von Wolfgang Stegmüller. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(1):138–151, March 1971. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170207.

Anonymous:1971:B

[43] Anonymous. [bibliographie]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(1):152–162, March 1971. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170208.

Anonymous:1971:VI

[44] Anonymous. Volume information. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(1):??, March 1971. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170194.

Kraft:1971:NBJ

[45] Victor Kraft. Nachruf auf Béla Juhos. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2 (2):163–173, September 1971. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170210.

Elzinga:1971:HTR

[46] Aant Elzinga. Huygens’ theory of research and Descartes’ theory of knowledge I. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(2):174–194, September 1971. CO-DEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170211.

Kisiel:1971:HNE

[47] Theodore Kisiel. Zu einer Hermeneutik naturwissenschaftlicher Entdeckung. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(2):195–221, September 1971. CO-DEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170212.

Schleichert:1971:PHG

[48] Hubert Schleichert. Zum Problem historischer Gesetze. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(2):222–238, September 1971. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170213.

Tornebohm:1971:FIS

[49] Håkan Törnebohm and Gerard Radnitzky. Forschung als innovatives System: Entwurf einer integrativen Sehweise, die Modelle erstellt zur Beschreibung und Kritik von Forschungsprozessen. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(2):239–290, September 1971. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170214.

Lauener:1971:WSG

[50] Henri Lauener. Wissenschaftstheorie in der Schweiz. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(2):291–317, September 1971. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170215.

Zaslawsky:1971:PSW

[51] Denis Zaslawsky. La philosophie des sciences (wissenschaftstheorie) en France (1950–1971). (French) [The philosophy of science in France (1950– 1971)]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(2):318–325, September 1971. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170216.

Juhos:1971:RKG

[52] Béla Juhos. Rezension: Konvergenzen in der Gegenwartsphilosophie und die moderne Physik von Walter v. Del-Negro. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(2):326–331, September 1971. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170217.

Freiberg:1971:RPW

[53] Christine Freiberg. Rezension: Philosophie und Wissenschaft. Eine Methodenlehre von Gerhard Frey. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(2):332, September 1971. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170218.

Essler:1971:RTG

[54] Wilhelm K. Essler. Rezension: Der Teil und das Ganze: Gespräche im Umkreis der Atomphysik von Werner Heisenberg. (German) [Book review: The Part and the Whole: Talks on the Field of Atomphysics by Werner Heisenberg]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(2):333–336, September 1971. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170219.

Kisiel:1971:REW

[55] Theodore Kisiel. Rezension: Einführung in die Wissenschaftstheorie von Helmut Seiffert. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(2):337, September 1971. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170220.

Anonymous:1971:NEB

[56] Anonymous. Nachträge und Ergänzungen zur Bibliographie der Schriften von Bela Juhos. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(2):338–339, September 1971. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170221.

Lauener:1971:VSW

[57] Henri Lauener. Veröffentlichungen schweizerischer Wissenschaftstheoretiker I. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(2):340–351, September 1971. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170222.

Anonymous:1971:BM

[58] Anonymous. Back matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(2):??, September 1971. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170223.

Bertman:1972:BPI

[59] Martin A. Bertman. Basic particulars and the identity thesis. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(1):1–8, March 1972. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170226.

Elzinga:1972:HTR

[60] Aant Elzinga. Huygens’ theory of research and Descartes’ theory of knowledge II. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(1):9–27, March 1972. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170227.

King-Farlow:1972:QTO

[61] John King-Farlow. Quantification theory and ontological monism: Logic as a tool for speculation. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(1):28–39, March 1972. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170228.

Sundaram:1972:KCS

[62] K. Sundaram. Kant or Cassirer: A study in complementarity. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(1):40–48, March 1972. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170229.

Ballestrem:1972:TBK

[63] K. Ballestrem and A. McCarthy. Thesen zur Begründung einer kritischen Theorie der Gesellschaft. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(1): 49–62, March 1972. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170230.

Beckermann:1972:RVK

[64] Ansgar Beckermann. Die realistischen Voraussetzungen der Konsenstheorie von J. Habermas. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(1):63–80, March 1972. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170231.

Rohs:1972:ZEB

[65] Peter Rohs. Ist die Zukunft eine Extrapolation? Bemerkungen zu einem Aufsatz von Victor Kraft. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(1):81– 84, March 1972. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170232.

Schmid:1972:FFG

[66] Michael Schmid. Falsifizierbarkeit oder Falsifikation?. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(1):85–87, March 1972. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170233.

Schnadelbach:1972:RNP

[67] Herbert Schnädelbach. Über den Realismus: Ein Nachtrag zum Positivismusstreit in der deutschen Soziologie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(1):88–112, March 1972. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170234.

Geldsetzer:1972:BIK

[68] Lutz Geldsetzer. Bericht über den 13. Internationalen Kongreß für Wissenschaftsgeschichte vom 18.–24. August 1971 in Moskau. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(1):113–118, March 1972. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170235.

Heidrich:1972:BIK

[69] Carl H. Heidrich. Bericht zum 4. Internationalen Kongreß für Logik, Methodologie und Philosophie der Wissenschaften vom 29. August bis 4. September 1971 in Bukarest. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(1): 119–126, March 1972. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170236.

Mirski:1972:WUG

[70] E. M. Mirski. Wissenschaftswissenschaft in der UdSSR (Geschichte, Probleme, Perspektiven). (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(1):127–144, March 1972. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170237.

Walentynowicz:1972:WWP

[71] Bohdan Walentynowicz. Wissenschaft der Wissenschaft in Polen. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(1):145–155, March 1972. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170238.

Kisiel:1972:RGH

[72] Theodore Kisiel. Rezension: Grundfragen der Hermeneutik. Ein philosophischer Beitrag (Philosophie in Einzeldarstellungen, 3. Band) von Emerich Coreth. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(1):156–158, March 1972. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170239.

Brown:1972:BRC

[73] Harold I. Brown. Book review: Criticism and the Growth of Knowledge by Imre Lakatos, Alan Musgrave. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(1):158–162, March 1972. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170240.

Schnadelbach:1972:REK

[74] Herbert Schnädelbach. Rezension: Das Elend der kritischen Theorie. Theodor W. Adorno. Herbert Marcuse. Jürgen Habermas von Günter Rohrmoser. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(1):162–166, March 1972. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170241.

Konig:1972:RNR

[75] Gert König. Rezension: Neue Reihen. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(1):167–169, March 1972. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170242.

Anonymous:1972:VWA

[76] Anonymous. Verzeichnis der wissenschaftlichen Arbeiten des korrespondierenden Mitglieds der Akademie der Wissenschaft der UdSSR P. W. Kopnin. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(1):170–175, March 1972. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170243.

Lauener:1972:VSW

[77] Henri Lauener. Veröffentlichungen schweizerischer Wissenschaftstheoretiker II. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(1):176–189, March 1972.

CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170244.

Anonymous:1972:BG

[78] Anonymous. [Bibliographie]. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(1):190– 200, March 1972. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170245.

Anonymous:1972:MHG

[79] Anonymous. Mitteilung der Herausgeber. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(1):201, March 1972. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170246.

Anonymous:1972:BM

[80] Anonymous. Back matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(1):??, March 1972. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170247.

Anonymous:1972:VI

[81] Anonymous. Volume information. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(1):??, March 1972. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170225.

Austin:1972:PRP

[82] William H. Austin. Paradigms, rationality, and partial communication. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(2):203–218, September 1972. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170249.

Bohler:1972:PAW

[83] Dietrich Böhler. Paradigmawechsel in analytischer Wissenschaftstheorie? Wissenschaftsgeschichtliche und Wissenschaftstheoretische Aufgaben der Philosophie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(2):219–242, September 1972. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170250.

Kluver:1972:WWE

[84] Jürgen Klüver and Wilfried Müller. Wissenschaftstheorie und Wissenschaftsgeschichte: Die Entdeckung der Benzolformel. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(2):243–266, September 1972. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170251.

Meyer:1972:WEU

[85] Willi Meyer. Wissenschaftstheorie und Erfahrung: Zur Überwindung des methodologischen Dogmatismus. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(2): 267–284, September 1972. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170252.

Prox:1972:SKG

[86] Lothar Prox. Strukturalistische Kunstforschung. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(2):285–297, September 1972. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170253.

Sagal:1972:ITN

[87] Paul T. Sagal. Incommensurability then and now. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3 (2):298–301, September 1972. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170254.

Becker:1972:DIH

[88] Werner Becker. Dialektik als Ideologie: Hegel und Marx: Eine kritische Betrachtung über Zustandekommen, Sinn und Funktion der dialektischen Methode. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(2):302–328, September 1972. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170255.

Keuth:1972:RPB

[89] Herbert Keuth. Eine Revolte in der Psychologie: Bemerkungen zu Holzkamps Vorschlag einer “kritisch-emanzipatorischen Psychologie”. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(2):329–357, September 1972. CO-DEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170256.

Kraft:1972:ZEG

[90] Victor Kraft. Ist die Zukunft eine Extrapolation?. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(2):358, September 1972. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170257.

Agazzi:1972:RDP

[91] Evandro Agazzi. Recent developments of the philosophy of science in Italy. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(2):359–371, September 1972. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170258.

vanPeursen:1972:WNG

[92] C. A. van Peursen and R. J. A. van Dijk. Wissenschaftstheorie in den Niederlanden. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(2):372–379, September 1972. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170259.

Brown:1972:BRJ

[93] Harold I. Brown. Book review: The Justification of Scientific Change by Carl R. Kordig. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(2):380–387, September 1972. CO-DEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170260.

Geldsetzer:1972:BRM

[94] Lutz Geldsetzer. Book review: Minnesota Studies in the Philosophy of Science (Herbert Feigl and Grover Maxwell, General Editors), Vol. IV: Analyses of Theories and Methods of Physics and Psychology by Michael Radner, Stephen Winokur. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(2):388–391, September 1972. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170261.

Anonymous:1972:KSW

[95] Anonymous. KongreßBerichte und Serien zur Wissenschaftstheorie I. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 3(2):392–400, September 1972. CO-DEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170262.

Agassi:1973:TBO

[96] Joseph Agassi. Testing as a bootstrap operation in physics. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(1):1–24, March 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170265.

Girill:1973:LSP

[97] T. R. Girill. The logic of scientific puzzles. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(1): 25–40, March 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170266.

Klowski:1973:UPL

[98] Joachim Klowski. Der unaufhebbare Primat der Logik, die Dialektik des Ganzen und die Grenze der Logik. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(1): 41–53, March 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170267.

Kraft:1973:GAG

[99] Victor Kraft. Die Giltigkeit von Aussagen. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(1):54–80, March 1973. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170268.

Stanage:1973:CHS

[100] Sherman M. Stanage. Cojectivity and the human sciences. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(1):81–97, March 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170269.

Munch:1973:RTE

[101] Richard Münch. Realismus und transzendentale Erkenntniskritik: Zur Kontroverse zwischen kritischem Rationalismus und Dialektik. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(1):98–107, March 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170270.

Silvers:1973:CTS

[102] Stuart Silvers. The critical theory of science. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(1): 108–132, March 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170271.

Bona:1973:LWU

[103] E. Bóna and J. Farkas. Die Lage der Wissenschaftstheorie in Ungarn. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(1):133–146, March 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170272.

Nordenstam:1973:PSN

[104] Tore Nordenstam and Hans Skjervheim. Philosophy of science in Norway. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(1):147–164, March 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170273.

Treml:1973:RAP

[105] Alfred K. Treml. Rezension: Analytische Philosophie I von W. K. Essler. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(1):165–169, March 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170274.

Kanitscheider:1973:RRR

[106] Bernulf Kanitscheider. Rezension: Realismus und Relativität von Joachim Hölling. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(1):170–174, March 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170275.

Kanitscheider:1973:RNA

[107] Bernulf Kanitscheider. Rezension: Neuzeit und Aufklärung von Jürgen Mittelstraß. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(1):174–178, March 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170276.

Heidrich:1973:BRT

[108] Carl H. Heidrich. Book review: Temporal Logic = Library of Exact Philosophy, Vol. 3 by Nicholas Rescher, Alasdair Urquhart. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(1):178–187, March 1973. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170277.

Bona:1973:VUW

[109] E. Bóna and J. Farkas. Veröffentlichungen ungarischer Wissenschaftstheoretiker. (German) [Publications of Hungarian philosophers of science]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(1):188–193, March 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170278.

Strosetzki:1973:BG

[110] Christoph Strosetzki. [Bibliographie]. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(1):194–201, March 1973. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170279.

Anonymous:1973:BM

[111] Anonymous. Back matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(1):??, March 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170280.

Anonymous:1973:VI

[112] Anonymous. Volume information. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(1):??, March 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170264.

Hammacher:1973:EMR

[113] Klaus Hammacher. Einige methodische Regeln Descartes’ und das erfindende Denken. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(2):203–223, September 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170282.

Kamlah:1973:IZG

[114] Andreas Kamlah. Invarianzgesetze und Zeitmetrik. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(2):224–260, September 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170283.

Kanitscheider:1973:GGP

[115] Bernulf Kanitscheider. Geochronometrie und Geometrodynamik: Zum Problem des Konventionalismus. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(2): 261–302, September 1973. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170284.

Klowski:1973:LSK

[116] Joachim Klowski. Läßt sich eine Kernlogik konstituieren? Ein Versuch dazu vom Standpunkt des pankritischen Rationalismus aus. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(2):303–312, September 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170285.

Kraft:1973:KEG

[117] Victor Kraft. Konstruktiver Empirismus. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(2):313–322, September 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170286.

Krimsky:1973:UMC

[118] Sheldon Krimsky. The use and misuse of critical gedankenexperimente. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(2):323–334, September 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170287.

Laszlo:1973:RGT

[119] Ervin Laszlo. The rise of general theories in contemporary science. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(2):335–344, September 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170288.

Frey:1973:RDL

[120] Gerhard Frey. Die Relevanz der deontischen Logik für die Ethik. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(2):345–355, September 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170289.

Weinberger:1973:ASA

[121] Ota Weinberger. Aufgaben und Schwierigkeiten der analytischen Rechtstheorie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(2):356–367, September 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170290.

Brandtstadter:1973:WGW

[122] Jochen Brandtstädter and Günther Reinert. Wissenschaft als Gegenstand der Wissenschaft vom menschlichen Erleben und Verhalten: Überlegungen zur Konzeption einer Wissenschaftspsychologie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(2):368–379, September 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170291.

Vollmer:1973:BRM

[123] Gerhard Vollmer. Book review: Method, Model and Matter by Mario Bunge. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(2):380–383, September 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170292.

Rapp:1973:REP

[124] Friedrich Rapp. Rezension: Erklärung. Prognose. Planung. Skizzen zu Brennpunktproblemen der Wissenschaftstheorie von Hans Lenk. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(2):383–385, September 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170293.

Schmid:1973:BRE

[125] Michael Schmid. Book review: Empiricism and the Analytic–Synthetic Distinction by Tore Nordenstam. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(2):385–388, September 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170294.

Kanitscheider:1973:BRO

[126] Bernulf Kanitscheider. Book review: Objective Knowledge by K. R. Popper. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(2):388–398, September 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170295.

Schmid:1973:BRH

[127] Michael Schmid. Book review: Human Understanding, Vol. I: The Collective Use and Evolution of Concepts by Stephen Toulmin. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(2):398–402, September 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170296.

Schmid:1973:BRL

[128] Michael Schmid. Book review: Law and Explanation. An Essay in the Philosophy of Science by Peter Achinstein. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(2): 402–407, September 1973. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170297.

Vollmer:1973:BRP

[129] Gerhard Vollmer. Book review: Philosophy of Physics by Mario Bunge. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 4(2):407–409, September 1973. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170298.

Embree:1974:NAG

[130] Lester Embree. Nachruf auf Aron Gurwitsch: Aron Gurwitsch als phänomenologischer Wissenschaftstheoretiker. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(1):1–8, March 1974. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170301.

Elzinga:1974:SRT

[131] Aant Elzinga. Some remarks on a theory of research in the work of Aristotle. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(1):9–38, March 1974. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170302.

Kamper:1974:HTP

[132] Dietmar Kamper. Hermeneutik: Theorie einer Praxis?. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(1):39–53, March 1974. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170303.

Kimmerle:1974:FHP

[133] Heinz Kimmerle. Die Funktion der Hermeneutik in den positiven Wissenschaften. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(1):54–73, March 1974. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170304.

Kramer:1974:OHW

[134] Hans Krämer. Zur Ortsbestimmung der historischen Wissenschaften. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(1):74–93, March 1974. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170305.

Tsouyopoulos:1974:IMI

[135] Nelly Tsouyopoulos. Die induktive Methode und das Induktionsproblem in der griechischen Philosophie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(1): 94–122, March 1974. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170306.

Heelan:1974:HES

[136] Patrick A. Heelan. Hermeneutics of experimental science in the context of the life-world. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(1):123–124, March 1974. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170307.

Kisiel:1974:CPH

[137] Theodore Kisiel. Commentary on Patrick Heelan’s “Hermeneutics of Experimental Science in the Context of the Life-World”. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(1):124–135, March 1974. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170308.

Heelan:1974:CPK

[138] Patrick A. Heelan. Comments on Professor Kisiel’s commentary. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(1):135–137, March 1974. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170309.

Kisiel:1974:NPS

[139] Theodore Kisiel and Galen Johnson. New philosophies of science in the USA: A selective survey. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(1):138–191, March 1974. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170310.

Braun:1974:RPP

[140] L. Braun. Rezension: Die Philosophie der Philosophiegeschichte im 19. Jahrhundert (Studien zur Wissenschaftstheorie Band III) von Lutz Geldsetzer. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(1):192–194, March 1974. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170311.

Geldsetzer:1974:RHD

[141] Lutz Geldsetzer. Rezension: Hermeneutik und Dialektik von R. Bubner, K. Cramer, R. Wiehl. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(1):194–196, March 1974. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170312.

Konig:1974:RLG

[142] Gert König. Rezension: Lexikon der Geschichte der Physik A–Z von Armin Hermann. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(1):197–200, March 1974. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170313.

Strosetzki:1974:BG

[143] Christoph Strosetzki. [Bibliographie]. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(1):201–209, March 1974. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170314.

Anonymous:1974:BM

[144] Anonymous. Back matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(1):??, March 1974. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170315.

Anonymous:1974:VI

[145] Anonymous. Volume information. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(1):??, March 1974. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170300.

Worrall:1974:NIL

[146] John Worrall. Nachruf auf Imre Lakatos: Imre Lakatos (1922–1974): Philosopher of Mathematics and Philosopher of Science. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(2):211–217, September 1974. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170317.

Buchel:1974:SWR

[147] Wolfgang Büchel. Die Struktur wissenschaftlicher Revolutionen und das Uhren-”Paradoxon”. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(2):218– 225, September 1974. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170318.

Detel:1974:ZFP

[148] Wolfgang Detel. Zwei Fallstudien zur Prüfung des Falsifikationismus. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(2):226–246, September 1974. CO-DEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170319.

Geldsetzer:1974:MMH

[149] Lutz Geldsetzer. Metaphilosophie als Metaphysik: Zur Hermeneutik der Bestimmung der Philosophie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(2):247– 255, September 1974. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170320.

Giesen:1974:REK

[150] Bernd Giesen and Michael Schmid. Rationalität und Erkenntnisfortschritt: Kritische Bemerkungen zur Methodologie der wissenschaftlichen Forschungsprogramme. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(2):256–284, September 1974. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170321.

Hubner:1974:FRF

[151] Kurt Hübner. Zur Frage des Relativismus und des Fortschritts in den Wissenschaften: Imre Lakatos zum Gedächtnis. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(2):285–303, September 1974. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170322.

Krah:1974:FWG

[152] Wolfgang Krah. Zum Falsifikationsprinzip in der Wissenschaftstheorie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(2):304–307, September 1974. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170323.

Ebert:1974:VEW

[153] Theodor Ebert. Über eine vermeintliche Entdeckung in der Wissenschaftstheorie: Anmerkungen zu Janich/Kambartel/Mittelstraß: Wissenschaftstheorie als Wissenschaftskritik. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(2): 308–316, September 1974. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170324.

Kraft:1974:GNG

[154] Victor Kraft. Die Giltigkeit von Normen. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(2):317–322, September 1974. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170325.

Porstmann:1974:WWA

[155] Reiner Porstmann. Werturteile als wissenschaftliche Aussagen? Zu Victor Krafts Beiträgen über die Giltigkeit von Aussagen und Normen. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(2):323–328, September 1974. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170326.

Rapp:1974:KEB

[156] Friedrich Rapp. Kybernetik und Erkenntnistheorie: Bemerkungen zur Konzeption von Georg Klaus. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(2): 329–340, September 1974. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170327.

Butts:1974:PSC

[157] Robert E. Butts. Philosophy of science in Canada. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5 (2):341–358, September 1974. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170328.

Geldsetzer:1974:BIK

[158] Lutz Geldsetzer. Bericht über den 14. Internationalen Kongreß für Wissenschaftsgeschichte vom 19.–27. August 1974 in Tokyo und Kyoto, Japan. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(2):359–362, September 1974. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170329.

Howson:1974:CSP

[159] Colin Howson and John Worrall. The contemporary state of philosophy of science in Britain. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(2):363–374, September 1974. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170330.

Tsouyopoulos:1974:RGA

[160] Nelly Tsouyopoulos. Rezension: Grundprobleme der antiken Wissenschaft von Kurt von Fritz. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(2):375–377, September 1974. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170331.

Schmid:1974:RGI

[161] Michael Schmid. Rezension: Gesellschaftstheorie und Ideologiekritik von Richard Münch. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(2):378–381, September 1974. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170332.

Rapp:1974:BRC

[162] Friedrich Rapp. Book review: The Coherence Theory of Truth by Nicholas Rescher. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(2):382–389, September 1974. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170333.

Butts:1974:VKW

[163] Robert E. Butts and John Galinaitis. Veröffentlichungen kanadischer Wissenschaftstheoretiker. (German) [Publications of Canadian philosophers of science]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(2):390–406, September 1974. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170334.

Anonymous:1974:EBG

[164] Anonymous. Eingegangene Bücher. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 5(2): 407–410, September 1974. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170335.

Frey:1975:LEN

[165] Gerhard Frey. Logik, Erfahrung und Norm: Zum Tode Victor Kraft’s. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(1):1–6, March 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170339.

Buchel:1975:SWS

[166] Wolfgang Büchel. Statistische Wahrscheinlichkeit und statistische Physik. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(1):7–18, March 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170340.

Keuth:1975:OPW

[167] Herbert Keuth. Objektivität und Parteilichkeit in der Wissenschaft. (german) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(1):19–33, March 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170341.

Rohatyn:1975:KHC

[168] Dennis A. Rohatyn. Kant, Hume and causality. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(1): 34–36, March 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170342.

Tornebohm:1975:EKF

[169] Håkan Törnebohm. An essay on knowledge-formation. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(1):37–64, March 1975. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170343.

Wandschneider:1975:EGU

[170] Dieter Wandschneider. Zur Eliminierung des Gödelschen Unvollständigkeitsproblems im Zusammenhang mit dem Antinomienproblem. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(1):65–81, March 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170344.

Freundlieb:1975:PDW

[171] Dieter Freundlieb. Zur Problematik einer Diskurstheorie der Wahrheit. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(1):82–107, March 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170345.

Harnatt:1975:BNS

[172] J. Harnatt. Zum Beweis der Nichtexistenz von “störenden Bedingungen”. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(1):108–112, March 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170346. See reply [231].

Heelan:1975:HRT

[173] Patrick A. Heelan. Heisenberg and radical theoretic change. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(1):113–136, March 1975. CODEN ZAWTA2.

ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://link.

springer.com/article/10.1007/BF01801105;   http://www.jstor.
org/stable/25170347.

Heisenberg:1975:CHT

[174] Werner Heisenberg and Patrick A. Heelan. Comments to Heelan’s thesis [with reply]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(1):137–138, March 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http:/

/link.springer.com/article/10.1007/BF01801106;   http://www.
jstor.org/stable/25170348.

Rapp:1975:MSB

[175] Friedrich Rapp. The methodological symmetry between verification and falsification. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(1):139–144, March 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170349.

Gochet:1975:RTP

[176] Paul Gochet. Recent trends in philosophy of science in Belgium. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(1):145–163, March 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170350.

Zecha:1975:RWH

[177] Gerhard Zecha. Rezension: Werturteilsstreit von Hans Albert, Ernst Topitsch. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(1):164–170, March 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170351.

Hess:1975:BRK

[178] Michael G. Hess. Book review: Kibernetika i metodologija nauki by B. V. Birjukov. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(1):170–176, March 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170352.

Morscher:1975:RBB

[179] Edgar Morscher. Rezension: Bernard Bolzanos Lehre von Anschauung und Begriff in ihrer Bedeutung für erkenntnistheoretische und pädagogische Probleme von Ursula Neemann. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(1): 176–178, March 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170353.

Frey:1975:NEB

[180] Gerhard Frey. Nachträge und Ergänzungen zur Bibliographie der Schriften von Victor Kraft. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(1):179–181, March 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170354.

Gochet:1975:PSB

[181] Paul Gochet. Philosophy of science in Belgium: A selective bibliography. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(1):182–186, March 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170355.

Seidel:1975:BG

[182] Inge Seidel and Dieter Zittlau. [Bibliographie]. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(1):187–202, March 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170356.

Anonymous:1975:FM

[183] Anonymous. Front matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(1):??, March 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170338.

Anonymous:1975:VI

[184] Anonymous. Volume information. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(1):??, March 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170337.

Braun:1975:EGF

[185] Günther E. Braun. Empirischer Gehalt und Falsifizierbarkeit: Eine semiotische Analyse des Popper-Kriteriums. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6 (2):203–216, September 1975. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170358.

Leist:1975:PBS

[186] Anton Leist. Ein Plädoyer für die Beendigung der Suche nach Wahrheitskriterien. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(2):217–234, September 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170359.

Weinberger:1975:FAG

[187] Ota Weinberger. Faktentranszendente Argumentation. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(2):235–251, September 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170360.

Wohlrapp:1975:AVK

[188] Harald Wohlrapp. Analytischer versus konstruktiver Wissenschaftsbegriff. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(2):252–275, September 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170361.

Ziemski:1975:TSR

[189] Stefan Ziemski. The typology of scientific research. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6 (2):276–291, September 1975. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170362.

Beckermann:1975:EBS

[190] Ansgar Beckermann. Einige Bemerkungen zur statistischen Kausalitätstheorie von P. Suppes. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(2):292– 310, September 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170363.

Braun:1975:SIK

[191] Günther E. Braun. ’Sollen impliziert Können’ und der entscheidungstheoretische Kontext: Zur Logik des Realisierbarkeitspostulates und seiner Anwendung. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(2):311–330, September 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170364.

Hoppe:1975:GSR

[192] Hansgeorg Hoppe. Goodmans Schein-Rätsel: Über die Widersprüchlichkeit und Erfahrungswidrigkeit des sog. “new riddle of induction”. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(2):331–339, September 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170365.

Stemmer:1975:GP

[193] Nathan Stemmer. The Goodman paradox. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(2):340– 354, September 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170366.

Dima:1975:PSR

[194] Teodor Dima. The philosophy of science in Romania. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6 (2):355–368, September 1975. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170367.

Schmid:1975:RAG

[195] Michael Schmid. Rezension: Analytische Geschichtsphilosophie von Karl Acham. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(2):369–371, September 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170368.

Dusberg:1975:RKP

[196] Klaus Jürgen Düsberg. Rezension: Konventionalität in der Physik von Werner Diederich. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(2):371–373, September 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170369.

Schmid:1975:BRS

[197] Michael Schmid. Book review: Sociology of Meaning by John O’Malley. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(2):374–377, September 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170370.

Buttemeyer:1975:RGI

[198] Wilhelm Büttemeyer. Rezension: Grundkurs im wissenschaftlichen Definieren. Übungen zum Selbststudium von Eike von Savigny. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(2):377–379, September 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170371.

Hegselmann:1975:RPE

[199] R. Hegselmann. Rezension: Pluralismus als Erkenntnismodell von Helmut Spinner. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(2):379–382, September 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170372.

Anonymous:1975:KSW

[200] Anonymous. KongreßBerichte und Serien zur Wissenschaftstheorie II. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(2):383–410, September 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170373.

Anonymous:1975:EBG

[201] Anonymous. Eingegangene Bücher. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(2): 411–412, September 1975. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170374.

Jammer:1976:WH

[202] Max Jammer. Werner Heisenberg (1901–1976). Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(1): 1–10, March 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170378.

Mercier:1976:FGM

[203] André Mercier. Ferdinand Gonseth (1890–1975): Mathematik und Wirklichkeit: Ein Stück Geschichte der Wissenschaftsphilosophie in der Schweiz. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(1):11–25, March 1976. CO-DEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170379.

Bourgeois:1976:VSS

[204] Warren Bourgeois. Verstehen in the social sciences. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(1):26–38, March 1976. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170380.

Gerber:1976:KKW

[205] Monika Gerber. Zur Korrespondenz- und Konsenstheorie der Wahrheit. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(1):39–57, March 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170381.

Hesse:1976:PBH

[206] Reinhard Hesse. Probleme der Begründungen von “Historische Größe” Ein Beitrag zur Kritik historischer Faktenkonstitution. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(1):58–74, March 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170382.

Kaufmann:1976:LSK

[207] Joseph N. Kaufmann. Löst der Strukturalismus die Kausalanalyse ab?. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(1):75–98, March 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170383.

Keuth:1976:PAP

[208] Herbert Keuth. Poppers Axiome für eine propensity-Theorie der Wahrscheinlichkeit. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(1):99–112, March 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170384.

Laucken:1976:VGE

[209] Uwe Laucken. Verstehen gegen Erklären: Nekrolog auf einen Gegensatz. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(1):113–118, March 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170385.

Eichner:1976:SZV

[210] Klaus Eichner. Zur Symmetrie zwischen Verifikation und Falsifikation. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(1):119–120, March 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170386.

Rapp:1976:HAR

[211] Friedrich Rapp. A helpful argument: Reply to K. Eichner. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(1):121–123, March 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170387.

Kleinknecht:1976:BVL

[212] Reinhard Kleinknecht. Bemerkungen über eine vermeintliche Lösung des Antinomienproblems. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(1):124– 126, March 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170388.

Pittioni:1976:KWK

[213] Veit Pittioni. Kritik an Wandschneiders Konstruktion. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(1):127–132, March 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170389.

Wandschneider:1976:RUA

[214] Dieter Wandschneider. Reflexive Unbeweisbarkeitsaussagen: Anmerkungen zur Grundsatz Diskussion um Gödels Unvollständigkeitsproblem. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(1):133–140, March 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170390.

Nowak:1976:SIM

[215] Leszek Nowak. On some interpretation of the Marxist methodology. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(1):141–183, March 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170391.

Schmid:1976:BRO

[216] Michael Schmid. Book review: Objectivity in Social Science by Frank Cunningham. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(1):184–186, March 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170392.

Schmid:1976:BRA

[217] Michael Schmid. Book review: Analytical Philosophy of Action by Arthur C. Danto. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(1):187–191, March 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170393.

Seiffert:1976:RSV

[218] Helmut Seiffert. Rezension: Von der Sprache zur Vernunft von Wilhelm Kamlah. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(1):191–192, March 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170394.

Seidel:1976:BG

[219] Inge Seidel and Dieter Zittlau. [Bibliographie]. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(1):193–208, March 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170395.

Anonymous:1976:FM

[220] Anonymous. Front matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(1):??, March 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170377.

Anonymous:1976:VI

[221] Anonymous. Volume information. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(1):??, March 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170376.

Bohme:1976:QMV

[222] Gernot Böhme. Quantifizierung — Metrisierung: Versuch einer Unterscheidung erkenntnistheoretischer und wissenschaftstheoretischer Momente im Prozeß der Bildung von quantitativen Begriffen. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(2):209–222, September 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170397.

Gavroglu:1976:RGP

[223] K. Gavroglu. Research guiding principles in modern physics: Case studies in elementary particle physics. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(2):223– 248, September 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170398.

Kanthack:1976:ZZP

[224] Lothar Kanthack and Ursula Wegener. Zum Zusammenhang zwischen Projektionsoperatoren und Eigenschaften. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(2):249–257, September 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170399.

Koppe:1976:HHD

[225] Franz Koppe. Die historisch-hermeneutischen Disziplinen im System der Wissenschaften. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(2):258–273, September 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170400.

Kuttner:1976:VDN

[226] Michael Küttner. Ein verbesserter deduktiv-nomologischer Erklärungsbegriff. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(2):274–297, September 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170401.

Pfarr:1976:PZR

[227] Joachim Pfarr. Die Protophysik der Zeit und das Relativitätsprinzip. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(2):298–326, September 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170402.

Weinryb:1976:HH

[228] Elazar Weinryb. Hermeneutics and history. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(2): 327–339, September 1976. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170403.

Ziemski:1976:TNM

[229] Stefan Ziemski. Towards a new model of science. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(2): 340–347, September 1976. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170404.

Braun:1976:KRP

[230] Günther E. Braun. Kritischer Rationalismus und politische Theorie: Zur Skizze eines sozialphilosophischen Forschungsprogramms. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(2):348–356, September 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170405.

Coffa:1976:RH

[231] J. Alberto Coffa. Reply to Harnatt. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(2):357– 358, September 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170406. See

[172].

Lorenzen:1976:EZG

[232] Paul Lorenzen. Die Eindeutigkeit der Zeitmessung. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(2):359–361, September 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170407.

Barbo:1976:GBU

[233] Francesca Barbò. Über Gödel’s Beweis der Unentscheidbarkeit: eine philosophische Bemerkung. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(2):362– 366, September 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170408.

Radnitzky:1976:PPF

[234] Gerard Radnitzky. Prinzipielle Problemstellungen der Forschungspolitik. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(2):367–403, September 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170409.

Schmid:1976:BRL

[235] Michael Schmid. Book review: The Logic of Society by Laird Addis. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(2):404–408, September 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170410.

Anonymous:1976:EBG

[236] Anonymous. Eingegangene Bücher. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(2): 409–411, September 1976. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170411.

Anonymous:1976:BM

[237] Anonymous. Back matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(2):??, September 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170412.

Poggeler:1977:MHG

[238] Otto Pöggeler. Martin Heidegger (1889–1976). (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(1):1–21, March 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170416.

Innis:1977:MMP

[239] Robert E. Innis. In memoriam Michael Polanyi (1891–1976). Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(1):22–29, March 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170417.

Agassi:1977:MRP

[240] Joseph Agassi. The methodology of research projects: A sketch. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(1):30–38, March 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170418.

Kaiser:1977:OGD

[241] Walter Kaiser. Operative Gesichtspunkte bei der Diskussion des Weberschen Gesetzes. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(1):39–47, March 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170419.

Pitt:1977:RSR

[242] Joseph C. Pitt and Morton Tavel. Revolutions in science and refinements in the analysis of causation. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(1):48–62, March 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170420.

Radermacher:1977:BTK

[243] Hans Radermacher. Der Begriff der Theorie in der kantischen und analytischen Philosophie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(1):63–76, March 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170421.

Schreiber:1977:IIL

[244] Alfred Schreiber. Das Induktionsproblem im Lichte der Approximationstheorie der Wahrheit. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(1):77–90, March 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25170422.

Backhaus:1977:POT

[245] Jürgen Backhaus. Politische Ökonomie als Theorie der Begründung: Eine Auseinandersetzung mit Friedrich Kambartels “Bemerkungen zum normativen Fundament der Ökonomie”. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(1): 91–117, March 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170423.

Kirchgassner:1977:NFO

[246] Gebhard Kirchgässner. Ein neues Fundament der Ökonomie?: Bemerkungen zu: Friedrich Kambartel, Bemerkungen zum normativen Fundament der Ökonomie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(1):118–131, March 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170424.

Kambartel:1977:USM

[247] Friedrich Kambartel. Zur Überwindung des Szientismus und Modellplatonismus in der Ökonomie: (Eine Erwiderung auf Jürgen Backhaus und Gebhard Kirchgäßner). (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(1):132– 143, March 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170425.

Sagal:1977:EEG

[248] Paul T. Sagal. Epistemology of Economics. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(1):144–162, March 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170426.

Kaminski:1977:DLP

[249] Stanisław Kaminski. The development of logic and the philosophy of science in Poland after the second world war. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(1): 163–171, March 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170427.

Firmenich:1977:RET

[250] Dieter Firmenich. Rezension: Entwurf einer Theorie des literarischen Gebildes von Horst–Jürgen Gerigk. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(1): 172–175, March 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170428.

Gethmann:1977:RMM

[251] Carl Friedrich Gethmann. Rezension: Marxismus als Methode. Wissenschaftstheoretische Untersuchungen zur Methode der marxistischen politischen Ökonomie von Christof Helberger. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(1):175–179, March 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170429.

Eecke:1977:BRR

[252] W. Ver Eecke. Book review: Rational Economic Man. A Philosophical Critique of Neo-Classical Economics by Martin Hollis, Edward Nell. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(1):179–182, March 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170430.

Seidel:1977:BG

[253] Inge Seidel and Dieter Zittlau. [Bibliographie]. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(1):183–193, March 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170431.

Anonymous:1977:FM

[254] Anonymous. Front matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(1):??, March 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170415.

Anonymous:1977:VI

[255] Anonymous. Volume information. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(1):??, March 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170414.

Aulin-Ahmavaara:1977:GTA

[256] A. Y. Aulin-Ahmavaara. A general theory of acts, with application to the distinction between rational and irrational ‘social cognition’. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(2):195–220, September 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170433.

Buchel:1977:BBF

[257] Wolfgang Büchel. Der Bellsche Beweis. Eine Fallstudie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(2):221–236, September 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170434.

Detel:1977:MEK

[258] Wolfgang Detel. Methode und Erkenntnisfortschritt: Kritische Bemerkungen zum Verhältnis von Wissenschaftstheorie und Wissenschaftsgeschichte. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(2):237–256, September 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170435.

Hoche:1977:KIB

[259] Hans-Ulrich Hoche. Kausalgefüge, irreale Bedingungssätze und das Problem der Definierbarkeit von Dispositionsprädikaten. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(2):257–291, September 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170436.

Hoyer:1977:ZNB

[260] Ulrich Hoyer. Ist das zweite Newtonsche Bewegungsaxiom ein Naturgesetz?. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(2):292–301, September 1977. CO-DEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170437.

Janich:1977:MTB

[261] Peter Janich. Ist Masse ein “theoretischer Begriff”?. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(2):302–314, September 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170438.

Thurnher:1977:FES

[262] Rainer Thurnher. Fall und Ereignis: Eine Studie zur Ontologie der Geschichtswissenschaft. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(2):315– 330, September 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170439.

Dusberg:1977:SWR

[263] Klaus Jürgen Düsberg. Stegmüller über “wissenschaftliche Revolutionen”. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(2):331–341, September 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170440.

Gethmann:1977:PBZ

[264] Carl Friedrich Gethmann and Rainer Hegselmann. Das Problem der Begründung zwischen Dezisionismus und Fundamentalismus. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(2):342–368, September 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170441.

Niessen:1977:DAW

[265] Manfred Niessen. Zur Differenz alltagsweltlicher und wissenschaftlicher Erklärungen. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(2):369–374, September 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170442.

Schuntermann:1977:AHS

[266] Michael F. Schuntermann. Zur Adäquatheit des Hacking–Stegmüllerschen Stützungsbegriffs. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(2):375– 378, September 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170443.

Wittrock:1977:FRM

[267] Björn Wittrock. Falsification, rejection, and modification. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(2):379–382, September 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170444.

Joerges:1977:WKE

[268] Bernward Joerges. Wissenschaftliche Kreativität Empirische und wissenschaftspraktische Hinweise. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(2): 383–404, September 1977. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170445.

Rapp:1977:RHB

[269] Friedrich Rapp. Rezension: In Honorem: Eine Bibliographie philosophischer Festschriften und ihrer Beiträge von Lutz Geldsetzer. (German) [].

Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(2):405–406, September 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170446.

Seiffert:1977:RPH

[270] Helmut Seiffert. Rezension: Die politische Herausforderung der Wissenschaft von Kurt Hübner, Nikolaus Lobkowicz, Hermann Lübbe, Gerard Radnitzky, Willy Hochkeppel. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(2): 406–409, September 1977. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170447.

Buttemeyer:1977:RGI

[271] Wilhelm Büttemeyer. Rezension: Grundkurs im logischen Schließen. Übungen zum Selbststudium von Eike von Savigny. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(2):409–410, September 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170448.

Anonymous:1977:EBG

[272] Anonymous. Eingegangene Bücher. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(2): 411–412, September 1977. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170449.

Anonymous:1977:BM

[273] Anonymous. Back matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 8(2):??, September 1977. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170450.

Schnelle:1978:YBH

[274] Helmut Schnelle. Yehoshua Bar-Hillel (1915–1975). Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(1):1–12, March 1978. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170454.

Lauener:1978:PB

[275] Henri Lauener. Paul Bernays (1888–1977). Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(1):

13–20, March 1978. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170455.

Grewendorf:1978:EAP

[276] Günther Grewendorf. Nicht-empirische Argumente: Zur Problematik ihrer wissenschaftstheorethischen Untersuchung. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(1):21–40, March 1978. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170456.

Kamlah:1978:MTU

[277] Andreas Kamlah. Metagesetze und theorie unabhängige Bedeutung physikalischer Begriffe. (German) [Meta laws and theory of independent meaning of physical terms]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(1):41–62, March 1978. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170457.

Lauener:1978:POG

[278] Henri Lauener. Probleme der Ontologie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(1):63–92, March 1978. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170458.

Stemmer:1978:RII

[279] Nathan Stemmer. The reliability of inductive inferences and our innate capacities. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(1):93–105, March 1978. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170459.

Strauss:1978:RKF

[280] Michael Strauss. Über Russells Kritik an Freges Unterscheidung zwischen Sinn und Bedeutung. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(1):106– 111, March 1978. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170460.

Ascheberg:1978:KPZ

[281] R. Ascheberg, G. Herrgott, and P. Krausser. Zur konstruktivistischen Protophysik der Zeit: Eine immanente Kritik. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(1):112–133, March 1978. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170461.

Raub:1978:BEM

[282] Werner Raub and Dirk Koppelberg. Bewertungen und Empfehlungen in der Methodologie wissenschaftlicher Forschungsprogramme. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(1):134–148, March 1978. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170462.

Wegener:1978:LWA

[283] Ursula Wegener. Eine Lösung der “wahrscheinlichkeitstheoretischen Antinomie der Quantenmechanik”. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(1): 149–156, March 1978. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170463.

Koppe:1978:LHIa

[284] Franz Koppe. Die literaturtheoretischen Hauptrichtungen und ihr Ertrag für eine Gegenstandsbestimmung der Literaturwissenschaft: Teil I: Positivismus und Formalismus. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(1): 157–184, March 1978. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170464.

Janssen:1978:RAS

[285] Paul Janssen. Rezension: Axiomatische Sprachtheorie. Wissenschaftstheoretische Untersuchungen zum Konstitutionsproblem der Einzelwissenschaften am Beispiel der Sprachwissenschaftstheorie Karl Bühlers von Rudolf Kamp. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(1):185–188, March 1978. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170465.

Gethmann-Siefert:1978:RWH

[286] Annemarie Gethmann-Siefert. Rezension: Wissenschaftstheorie. Handlungstheorie. Fundamentale Theologie von Helmut Peukert. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(1):188–192, March 1978. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170466.

Seidel:1978:BG

[287] Inge Seidel and Dieter Zittlau. [Bibliographie]. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(1):193–205, March 1978. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170467.

Anonymous:1978:BM

[288] Anonymous. Back matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(1):??, March 1978. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170468.

Anonymous:1978:FM

[289] Anonymous. Front matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(1):??, March 1978. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170453.

Anonymous:1978:VI

[290] Anonymous. Volume information. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(1):??, March 1978. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170452.

Bratoev:1978:BEK

[291] Georgy Bratoev. Bemerkungen über die Explikation der kausalen Begriffe. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(2):207–224, September 1978. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170470.

Engels:1978:TSF

[292] Eve-Marie Engels. Teleologie: eine “Sache der Formulierung” oder eine “Formulierung der Sache”?: Überlegungen zu Ernest Nagels reduktionistischer Strategie und Versuch ihrer Widerlegung. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(2):225–235, September 1978. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170471.

Keuth:1978:MRK

[293] Herbert Keuth. Methodologische Regeln des kritischen Rationalismus Eine Kritik. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(2):236–255, September 1978. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170472.

Lorenzen:1978:KDD

[294] Paul Lorenzen. Eine konstruktive Deutung des Dualismus in der Wahrscheinlichkeitstheorie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(2):256– 275, September 1978. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170473.

Lugg:1978:DS

[295] Andrew Lugg. Disagreement in science. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(2):276– 292, September 1978. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170474.

Steinvorth:1978:WWM

[296] Ulrich Steinvorth. Wertfreiheit der Wissenschaften bei Marx, Weber und Adorno: Ein Nachtrag zum Methodenstreit zwischen Kritischer Theorie und Kritischem Rationalismus. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(2): 293–306, September 1978. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170475.

Trapp:1978:EPG

[297] Rainer W. Trapp. Exaktheit in der Philosophie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(2):307–336, September 1978. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170476.

Eichner:1978:NCB

[298] Klaus Eichner and Werner Habermehl. A note on the controversy between Rapp and Eichner. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(2):337–338, September 1978. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170477.

Rapp:1978:DCA

[299] Friedrich Rapp. Disentangling counter-arguments. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9 (2):339–342, September 1978. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170478.

Janich:1978:PZR

[300] Peter Janich. Die Protophysik der Zeit und das Relativitätsprinzip: Erwiderung auf Joachim Pfarr. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(2):343– 347, September 1978. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170479.

Tietzel:1978:FVL

[301] Manfred Tietzel. Die Finalisierungsdebatte oder: Viel Lärm um nichts. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(2):348–360, September 1978. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170480.

Koppe:1978:LHIb

[302] Franz Koppe. Die literaturtheoretischen Hauptrichtungen und ihr Ertrag für eine Gegenstandsbestimmung der Literaturwissenschaft: Teil II: Strukturalismus. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(2):361–398, September 1978. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170481.

Geldsetzer:1978:RSG

[303] Lutz Geldsetzer. Rezension: Seminar: Geschichte und Theorie. Umrisse einer Historik von Hans-Michael Baumgartner, Jörn Rüsen. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(2):399–404, September 1978. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170482.

Rehder:1978:RWM

[304] Wulf Rehder. Rezension: Wider den Methodenzwang. Skizze einer anarchistischen Erkenntnistheorie von Paul Feyerabend. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(2):404–413, September 1978. CODEN ZAWTA2.

ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170483.

Pahler:1978:BRP

[305] Klaus Pähler. Book review: The Philosophy of Karl Popper by P. A. Schilpp. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(2):413–422, September 1978. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170484.

Braun:1978:RWW

[306] Günther E. Braun. Rezension: Zur Wissenschaftstheorie der Wirtschaftsplanung von Frithjof Spreer. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(2):423– 425, September 1978. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170485.

Anonymous:1978:EBG

[307] Anonymous. Eingegangene Bücher. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 9(2): 426–430, September 1978. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170486.

Boon:1979:RTR

[308] Louis Boon. Repeated tests and repeated testing: How to corroborate low level hypotheses. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(1):1–10, March 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170490.

Dusberg:1979:ETF

[309] Klaus Jürgen Düsberg. Sind empirische Theorien falsifizierbar?. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(1):11–27, March 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170491.

Hoyer:1979:VLK

[310] Ulrich Hoyer. Das Verhältnis der Leibnizschen zur Keplerschen Himmelsmechanik. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(1):28–34, March 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170492.

Meran:1979:IKV

[311] Josef Meran. Individualismus oder Kollektivismus?: Versuch einer Rekonstruktion eines sozialwissenschaftlichen Grundlagenstreits. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(1):35–53, March 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170493.

Nordenstam:1979:RED

[312] Tore Nordenstam and Håkan Törnebohm. Research, ethics and development. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(1):54–66, March 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170494.

Radnitzky:1979:PTB

[313] Gerard Radnitzky. Das Problem der Theorienbewertung: Begründungsphilosophischer, skeptischer und fallibilistischer Denkstil in der Wissenschaftstheorie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(1):67–97, March 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170495.

Schock:1979:CH

[314] Rolf Schock. On classifications and hierarchies. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(1): 98–106, March 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170496.

Shmueli:1979:HOS

[315] Efraim Shmueli. How is objectivity in the social sciences possible?: A re-evaluation of Karl Mannheim’s concept of “relationism”. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(1):107–118, March 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170497.

Brown:1979:FAS

[316] Harold I. Brown. A functional analysis of scientific theories. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(1):119–140, March 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170498.

Stehr:1979:DMD

[317] Nico Stehr and Anthony Simmons. The Diversity of Modes of Discourse and the Development of Sociological Knowledge. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(1):141–161, March 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170499.

Danneberg:1979:VLA

[318] Lutz Danneberg and Hans-Harald Müller. Verwissenschaftlichung der Literaturwissenschaft Ansprüche, Strategien, Resultate. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(1):162–191, March 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170500.

Witzleben:1979:RHS

[319] Frank Witzleben. Rezension: Handlung, Sprache und Vernunft von Rüdiger Bubner. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(1):192–196, March 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170501.

Lorenzen:1979:RMW

[320] Paul Lorenzen. Rezension: Mathematik und Weltanschauung von Wolfram Heitsch. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(1):197–200, March 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170502.

Hayakawa:1979:RLP

[321] Jiro Hayakawa. Rezension: Logik der philosophischen Erkenntnis von Shiro Nagai. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(1):200–201, March 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170503.

Seidel:1979:BG

[322] Inge Seidel and Dieter Zittlau. [Bibliographie]. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(1):202–211, March 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170504.

Anonymous:1979:BMG

[323] Anonymous. Back Matter. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(1):??, March 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25170505.

Anonymous:1979:FM

[324] Anonymous. Front matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(1):??, March 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170489.

Anonymous:1979:VI

[325] Anonymous. Volume information. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(1):??, March 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170488.

Baldamus:1979:EPW

[326] W. Baldamus. Das exoterische Paradox der Wissenschaftsforschung: Ein Beitrag zur Wissenschaftstheorie Ludwik Flecks. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(2):213–233, September 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170507.

Braun:1979:THP

[327] Lucien Braun. Théorie et histoire de la philosophie. (French) [Theory and history of philosophy]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(2):234–243, September 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170508.

Cohen:1979:NVG

[328] Yaël Cohen. A new view of Grue. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(2):244– 252, September 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170509.

Klein:1979:FMT

[329] Jürgen Klein. On the function and meaning of theory, intellectual and social history for a reconstruction of German “anglistik” as a rational discipline. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(2):253–266, September 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170510.

Riedel:1979:UEW

[330] Manfred Riedel. Die Universalität der europäischen Wissenschaft als begriffs- und wissenschaftsgeschichtliches Problem. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(2):267–287, September 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170511.

Rodi:1979:DPG

[331] Frithjof Rodi. Diesseits der Pragmatik: Gedanken zu einer Funktionsbestimmung der hermeneutischen Wissenschaften. (German) [on this side of pragmatics: Thoughts on a function determining the hermeneutic sciences ]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(2):288–315, September 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170512.

Sass:1979:EWK

[332] Hans-Martin Sass and A. Einstein. Einstein über “wahre Kultur” und die Stellung der Geometrie im Wissenschaftssystem: Ein Brief Albert Einsteins an Hans Vaihinger vom Jahr 1919. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(2):316–319, September 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170513.

Tietzel:1979:WCG

[333] Manfred Tietzel. War Cäsar ein Großer Mann?. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(2):320–337, September 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170514.

Ziemski:1979:TTS

[334] Stefan Ziemski. Two Types of Scientific Research. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(2):338–342, September 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170515.

Bohme:1979:WKA

[335] Gernot Böhme. Wie kann es abgeschlossene Theorien geben?. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(2):343–351, September 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170516.

Buchel:1979:FWL

[336] Wolfgang Büchel. Finalisierung der Wissenschaft und Lyssenkoismus. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(2):352–357, September 1979. CO-DEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170517.

Dusberg:1979:BBA

[337] Klaus Jürgen Düsberg. Bemerkung zum Begriff der “abgeschlossenen” Theorie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(2):358–362, September 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170518.

Finocchiaro:1979:MCC

[338] Maurice A. Finocchiaro. Methodological criticism and critical methodology: An analysis of Popper’s critique of Marxian social science. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(2):363–374, September 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170519.

Keuth:1979:EEK

[339] Herbert Keuth. Erkenntnis oder Entscheidung? Die Konsenstheorien der Wahrheit und der Richtigkeit von Jürgen Habermas. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(2):375–393, September 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170520.

Geldsetzer:1979:NTP

[340] Lutz Geldsetzer. Neueste Tendenzen und Problemstellungen der Philosophie und Wissenschaftstheorie in der Bundesrepublik Deutschland aus der Sicht sowjetischer Beobachter. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(2):

394–404, September 1979. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170521.

Vollmer:1979:BRF

[341] Gerhard Vollmer. Book review: The Furniture of the World by M. Bunge. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(2):405–407, September 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170522.

Koppelberg:1979:RWR

[342] Dirk Koppelberg. Rezensionen: Der wissenschaftstheoretische Realismus und die Autorität der Wissenschaften von Paul Feyerabend; Erkenntnis für freie Menschen von Paul Feyerabend. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(2):407–414, September 1979. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170523.

Anonymous:1979:EBG

[343] Anonymous. Eingegangene Bücher. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 10(2): 414–416, September 1979. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170524.

Hammer:1980:PWI

[344] Felix Hammer. Perspektiven einer Wissenschaftsethik im Dialog mit Francis Bacon. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(1):1–15, March 1980. CO-DEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170528.

Johansson:1980:CPC

[345] Ingvar Johansson. Ceteris paribus clauses, closure clauses and falsifiability. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(1):16–22, March 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170529.

Kamlah:1980:WAA

[346] Andreas Kamlah. Wie arbeitet die analytische Wissenschaftstheorie?. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(1):23–44, March 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170530.

Sass:1980:QHS

[347] Hans-Martin Sass. The quest for humanism in a scientific society. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(1):45–52, March 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170531.

Stachowiak:1980:MEG

[348] Herbert Stachowiak. Der Modellbegriff in der Erkenntnistheorie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(1):53–68, March 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170532.

Strube:1980:PAS

[349] Werner Strube. Philosophische Analyse der Sprache sprachanalytischer Philosophen. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(1):69–79, March 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170533.

Wegener:1980:LCP

[350] Ursula Wegener. Sind Ludwigs Chancengewichtungen propensities im Sinne Poppers?. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(1):80–85, March 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170534.

Agassi:1980:SS

[351] Joseph Agassi and John R. Wettersten. Stegmüller squared. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(1):86–94, March 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170535.

Charpa:1980:KEG

[352] Ulrich Charpa. Konstruktivistische Erklärungstheorie als Grundlage einer methodisch verfahrenden Philologie?. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(1):95–102, March 1980. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170536.

Dusberg:1980:KND

[353] Klaus Jürgen Düsberg. Zur Kritik einer neuen Definition von “Wahrscheinlichkeit”. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(1):103– 107, March 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170537.

Kutschmann:1980:MEW

[354] Werner Kutschmann. Zu “Methode und Erkenntnisfortschritt” von Wolfgang Detel. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(1):108–114, March 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170538.

Niemeyer:1980:DMA

[355] Christian Niemeyer. “Der Drache und das Miezekätzchen”: Anmerkungen zu einer Wissenschaftskonzeption, die ein Märchen bleibt oder: Was bringt ein Vergleich zwischen Feyerabend und Holzkamp?. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(1):115–134, March 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170539.

Steinvorth:1980:LPT

[356] Ulrich Steinvorth. Lakatos und politische Theorie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(1):135–146, March 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170540.

Weinberger:1980:WAK

[357] Christiane Weinberger. Wissenschaftstheoretische Anmerkungen zu Konrad Lorenz’ “Vergleichende Verhaltensforschung”. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(1):147–161, March 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170541.

Merkens:1980:BPD

[358] Hans Merkens. Bemerkung zur Popper-Diskussion. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(1):162–163, March 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170542.

Poggeler:1980:TRP

[359] Otto Pöggeler. Is there research policy making vis-à-vis the geisteswissenschaften? Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(1):164–193, March 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170543.

Geldsetzer:1980:BRL

[360] Lutz Geldsetzer. Book review: Logic, Foundations of Mathematics and Computability Theory/Foundational Problems in the Special Sciences/ Basic Problems in Methodology and Linguistics/Historical and Philosophical Dimensions of Logic, Methodology and Philosophy of Science. Parts One, Two, Three and Four of the Proceedings of the Fifth International Congress of Logic, Methodology and Philosophy of Science by R. E. Butts, J. Hintikka. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(1):194–195, March 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170544.

Charpa:1980:RWL

[361] Ulrich Charpa. Rezension: Wissenschaftstheoretisches Lexikon von E. Braun, H. Radermacher. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(1):196– 199, March 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170545.

Coersmeier:1980:BG

[362] Birgit Coersmeier and Larry Steindler. [Bibliographie]. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(1):200–212, March 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170546.

Anonymous:1980:FM

[363] Anonymous. Front matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(1):??, March 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170527.

Anonymous:1980:VI

[364] Anonymous. Volume information. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(1):??, March 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170526.

Baltas:1980:MPT

[365] A. Baltas and K. Gavroglu. A modification of Popper’s tetradic schema and the special relativity theory. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(2):213–237, September 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170548.

Dougherty:1980:BGP

[366] F. W. P. Dougherty. Buffon’s Gnoseological Principle. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(2):238–253, September 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170549.

Tietzel:1980:SDL

[367] Manfred Tietzel. Sympathy for the Devil: Die literarische figur des wissenschaftlers aus der sicht der modernen wissenschaftslehre. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(2):254–275, September 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170550.

Tuomela:1980:AD

[368] Raimo Tuomela. Analogy and distance. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(2):276– 291, September 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170551.

Zahar:1980:PKG

[369] Elie Zahar. Positivismus und Konventionalismus. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(2):292–301, September 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170552.

Neemann:1980:EKT

[370] Ursula Neemann. Ist Eckard Königs “Theorie der Erziehungswissenschaft” eine “Kritik der pädagogischen Vernunft”?. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(2):302–320, September 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170553.

Krah:1980:FNT

[371] Wolfgang Krah. Zur Forderung nach theoretischem Pluralismus in Permanenz. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(2):321–331, September 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170554.

Rolf:1980:BHV

[372] Bertil Rolf. Black and Hempel on vagueness. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(2): 332–346, September 1980. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170555.

Schopman:1980:FF

[373] J. Schopman. Finalization and functionalization. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(2):347–353, September 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170556.

Schuntermann:1980:HLL

[374] Michael F. Schuntermann. Hackings Law of Likelihood und das Kartenparadoxon von Kerridge. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(2):354– 356, September 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170557.

Wegener:1980:VLB

[375] Ursula Wegener. Ein Vergleich der von Ludwig bzw. Popper vorgeschlagenen Interpretationen der Quantenmechanik. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(2):357–366, September 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170558.

Weismann:1980:SQF

[376] Annabella Weismann. Sozialforschung und Quantifizierung: Ein forschungsstrategischer Beitrag zur Wissenschaftsgeschichte der empirischen Sozialforschung. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(2):367– 384, September 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170559.

Firmenich:1980:RWL

[377] Dieter Firmenich. Rezensionen: Zur Wissenschaftstheorie der Literaturwissenschaft von Dieter Freundlieb; Argumentation in der Literaturwissenschaft von Eike von Savigny; Die Sprache der Literaturwissenschaft von Harald Fricke. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(2):385–387, September 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170560.

Welter:1980:RSB

[378] Rüdiger Welter. Rezension: Sprache und Bedürfnis. Zur sprachphilosophischen Grundlage der Geisteswissenschaften von Franz Koppe. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(2):387–393, September 1980. CO-DEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170561.

Dontschev:1980:RPW

[379] Gentscho Dontschev. Rezension: Probleme der wissenschaftlichen Erkenntnis vom methodologischen Gesichtspunkt aus von Azari Polikarov. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(2):393–397, September 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170562.

Schmid:1980:RFR

[380] Michael Schmid. Rezension: Fortschritt und Rationalität der Wissenschaft von Gerhard Radnitzky, Gunnar Anderson. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(2):397–405, September 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170563.

Buttemeyer:1980:RKN

[381] Wilhelm Büttemeyer. Rezension: Klassische und nichtklassische Aussagenlogik von Wolfgang Rautenberg. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(2):405–407, September 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170564.

Zittlau:1980:BRH

[382] Dieter Zittlau. Book review: Hans Reichenbach, Selected Writings, 1909– 1953 by Maria Reichenbach, Robert S. Cohen; Hans Reichenbach, Logical Empiricist by Wesley C. Salmon. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(2):407– 412, September 1980. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170565.

Geldsetzer:1980:BRP

[383] Lutz Geldsetzer. Book review: Philosophy of History and Action by Yirmiahu Yovel. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(2):412–414, September 1980. CO-DEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170566.

Anonymous:1980:EBG

[384] Anonymous. Eingegangene Bücher. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 11(2): 414–418, September 1980. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170567.

Marbach:1981:JP

[385] Eduard Marbach. Jean Piaget (1896–1980). Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(1): 1–27, March 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170571.

Hoche:1981:MKA

[386] Hans-Ulrich Hoche. Zur Methodologie von Kombinationstests in der analytischen Philosophie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(1):28– 54, March 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170572.

Ludwig:1981:ABP

[387] Günther Ludwig. Axiomatische Basis einer physikalischen Theorie und theoretische Begriffe. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(1):55– 74, March 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170573.

Neemann:1981:ULF

[388] Ursula Neemann. Zur Unterscheidung von logischer und faktischer Wahrheit. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(1):75–97, March 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170574.

Pahler:1981:TEB

[389] Klaus Pähler. Teststrenge und empirische Bewährung in der Popperianischen Wissenschaftstheorie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(1):98– 109, March 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170575.

Schmidt:1981:REL

[390] Klaus J. Schmidt. Ein Rechenverfahren für die elementare Logik. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(1):110–115, March 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170576.

Busshoff:1981:WPT

[391] Heinrich Busshoff. Wissenschaftstheorie und politische Theorie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(1):116–134, March 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170577.

Charpa:1981:AAK

[392] Ulrich Charpa. Eine Anmerkung zu Andreas Kamlahs Darstellung der ‘normativ-analytischen’ Wissenschaftstheorie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(1):135–137, March 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170578.

Kamlah:1981:MDA

[393] Andreas Kamlah. Methode oder Dogma?: Eine Auseinandersetzung mit der zweiten Auflage von P. Janichs Protophysik der Zeit. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(1):138–162, March 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170579.

Kuttner:1981:TNS

[394] Michael Küttner. Theorie unter dem Non-Statement View und der Kuhnsche Wissenschaftler. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(1):163– 177, March 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170580.

Vollmer:1981:BRW

[395] Gerhard Vollmer. Book review: A World of Systems by M. Bunge. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(1):178–179, March 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170581.

Mainzer:1981:RVE

[396] Klaus Mainzer. Rezension: Vernuft, Erkenntnis, Sittlichkeit (Intern. Philos. Symposium, Göttingen vom 27.–29. Oktober 1977 aus Anlaß d. 50. Todestages von Leonard Nelson) Hamburg 1979. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(1):180–182, March 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170582.

Kamlah:1981:RMR

[397] Andreas Kamlah. Rezension: Zur Messung von Raum und Zeit. Eine Kritik der sogenannten Protophysik von K. J. Düsberg. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(1):182–187, March 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170583.

Coersmeier:1981:BG

[398] Birgit Coersmeier and Larry Steindler. [Bibliographie]. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(1):188–207, March 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170584.

Anonymous:1981:BM

[399] Anonymous. Back matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(1):??, March 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170585.

Anonymous:1981:FM

[400] Anonymous. Front matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(1):??, March 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170570.

Anonymous:1981:VI

[401] Anonymous. Volume information. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(1):??, March 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170569.

Carrier:1981:GFI

[402] Martin Carrier. Goethes Farbenlehre: ihre Physik und Philosophie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(2):209–225, September 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170587.

Dascal:1981:PSR

[403] Marcelo Dascal and Asher Idan. Procedures in scientific research and in language understanding. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(2):226–249, September 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170588.

Jokisch:1981:KIU

[404] Rodrigo Jokisch. Kraft und Identität: Überlegungen zur anthropologischen Voraussetzung der neuzeitlichen Wissenschaft. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(2):250–262, September 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170589.

McQueen:1981:SSE

[405] David V. McQueen. Sociology as a science: An essay on the roots of the “scientific method” and its relationship to the pursuit of sociology. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(2):263–284, September 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170590.

Schock:1981:ITR

[406] Rolf Schock. The inconsistency of the theory of relativity. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(2):285–296, September 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170591.

Schopman:1981:HSE

[407] Joop Schopman. The history of semiconductor electronics: A Kuhnian story? Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(2):297–302, September 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170592.

Strube:1981:SPE

[408] Werner Strube. Ist die sprachanalytische Philosophie einzelsprachlich relativ?. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(2):303–321, September 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170593.

Audretsch:1981:QGS

[409] Jürgen Audretsch. Quantum gravity and the structure of scientific revolutions. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(2):322–339, September 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170594.

Muhlholzer:1981:PZE

[410] Felix Mühlhölzer. Zur Protophysik der Zeit: Eine erneute Kritik. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(2):340–352, September 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170595.

Regelmann:1981:ALH

[411] Johann-Peter Regelmann. Die Aktualität Lyssenkos: Historische Ergänzungen zu einer wissenschaftstheoretischen Debatte. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(2):353–363, September 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170596.

Koppelberg:1981:EWA

[412] Dirk Koppelberg. Ende oder Wende der analytischen Philosophie und Wissenschaftstheorie?: Einige Bemerkungen zum Hegel-Kongreß 1981 in Stuttgart: “Kant oder Hegel? Über Formen der Begründung in der Philosophie”. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(2):364–400, September 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170597.

Braun:1981:RMF

[413] Hans-Joachim Braun. Rezension: Die Macht des Fortschritts. Plädoyer für Technik und Wissenschaft von Wolfgang Büchel. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(2):401–403, September 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170598.

Kroczek:1981:RWV

[414] Ronald Kroczek. Rezension: Wille, Vorstellung und Wirklichkeit. Menschliche Erkenntnis und physikalische Naturbeschreibung von Wolfgang Seelig. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(2):403–404, September 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170599.

Brendgen:1981:RSF

[415] Karl-Heinz Brendgen. Rezension: Die Struktur und Funktion transzendentaler Argumentationsfiguren. Ein argumentationstheoretischer Beitrag zur Wissenschaftsphilosophie von Peter E. Stüben. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(2):404–411, September 1981. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170600.

Anonymous:1981:EBG

[416] Anonymous. Eingegangene Bücher. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 12(2): 411–412, September 1981. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170601.

Schramm:1982:WH

[417] Matthias Schramm. Willy Hartner (1905–1981). Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(1): 1–21, March 1982. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170605.

Balzer:1982:KSG

[418] Wolfgang Balzer and Felix Mühlhölzer. Klassische Stoßmechanik. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(1):22–39, March 1982. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170606.

Buchel:1982:TNG

[419] Wolfgang Büchel. Teleologie und Negentropie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(1):40–47, March 1982. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170607.

Steinvorth:1982:MWS

[420] Ulrich Steinvorth. Max Webers System der verstehenden Soziologie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(1):48–69, March 1982. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170608.

Tetens:1982:WNG

[421] Holm Tetens. Was ist ein Naturgesetz?. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(1):70–83, March 1982. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170609.

Will:1982:PRI

[422] Ulrich Will. Eine pragmatische Rechtfertigung der Induktion. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(1):84–98, March 1982. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170610.

Radder:1982:ICL

[423] Hans Radder. An immanent criticism of Lakatos’ account of the ‘degenerating phase’ of Bohr’s atomic theory. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(1):99– 109, March 1982. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170611. See re-

sponse [465].

Radnitzky:1982:KGI

[424] Gerard Radnitzky. Knowing and guessing: If all knowledge is conjectural, can we then speak of cognitive progress? On persistent misreadings of Popper’s work. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(1):110–121, March 1982. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170612.

Engels:1982:TTU

[425] Eve-Marie Engels. Teleologie ohne Telos? Überlegungen zum XIX. Symposium der Gesellschaft für Wissenschaftsgeschichte e.V. vom 28.–30. Mai 1981 in Bamberg über “Die Idee der Zweckmäßigkeit in der Geschichte der Wissenschaften”. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(1):122–165, March 1982. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170613.

vanThemaat:1982:BRL

[426] Verloren van Themaat. Book review: Laboratory Life. The Social Construction of Scientific Facts by Bruno Latour, Steve Woolgar. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(1):166–170, March 1982. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170614.

Wettersten:1982:BRM

[427] John R. Wettersten. Book review: Models. Representations and the Scientific Understanding by Marx W. Wartofsky. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(1): 170–173, March 1982. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170615.

Schramm:1982:BWH

[428] Matthias Schramm. Bibliographie Willy Hartner. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(1):174–179, March 1982. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170616.

Coersmeier:1982:BG

[429] Birgit Coersmeier and Larry Steindler. [Bibliographie]. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(1):180–206, March 1982. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170617.

Anonymous:1982:BM

[430] Anonymous. Back matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(1):??, March 1982. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170618.

Anonymous:1982:FM

[431] Anonymous. Front matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(1):??, March 1982. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170604.

Anonymous:1982:VI

[432] Anonymous. Volume information. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(1):??, March 1982. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170603.

Bunge:1982:CBP

[433] Mario Bunge. Is chemistry a branch of physics? Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(2): 209–223, September 1982. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170620.

Carter:1982:GDS

[434] Richard B. Carter. Gilbert and Descartes: The science of conserving the compound body. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(2):224–233, September 1982.

CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170621.

Jaenecke:1982:GMG

[435] Peter Jaenecke. Grundzüge einer Meßtheorie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(2):234–279, September 1982. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170622.

Knapp:1982:ZWE

[436] Hans G. Knapp. Zwei Wege der Erkenntnis: Ein Beitrag zur Rekonstruktion von Erkenntnisprozessen. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(2): 280–293, September 1982. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170623.

Lorenz:1982:RFL

[437] Dieter Lorenz. Über die “Realität” der FitzGerald–Lorentz-Kontraktion. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(2):294–319, September 1982. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170624.

Ros:1982:KTT

[438] Arno Ros. Kausale, teleologische und teleonomische Erklärungen. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(2):320–335, September 1982. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170625.

Weinert:1982:AGV

[439] Friedel Weinert. Die Arbeit der Geschichte: Ein Vergleich der Analysemodelle von Kuhn und Foucault. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(2): 336–358, September 1982. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170626.

Rasmussen:1982:FCT

[440] Stig Alstrup Rasmussen. Finalization and completed theories. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(2):359–369, September 1982. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170627.

Steinvorth:1982:BWP

[441] Ulrich Steinvorth. Bußhoffs Wissenschaftstheorie der Politikwissenschaft und Lakatos. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(2):370–382, September 1982. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170628.

Friedmann:1982:BLI

[442] Johannes Friedmann. Bemerkungen zur Logikbegründung im Deutschen Konstruktivismus. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(2):383–402, September 1982. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170629.

Jacobi:1982:REM

[443] Klaus Jacobi. Rezension: Einführung in die Methodologie. Elementare allgemeine wissenschaftliche Denkmethoden im Überblick von Albert Menne. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(2):403–412, September 1982. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170630.

Anonymous:1982:EBG

[444] Anonymous. Eingegangene Bücher. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 13(2): 412–414, September 1982. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170631.

Gross:1983:GST

[445] Helmut Gross. Grundsatzfragen sozialwissenschaftlicher Theoriebildung. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 14(1):1–14, March 1983. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170635.

Klein:1983:OTI

[446] Jürgen Klein, Vera Damm, and Angelika Giebeler. An outline of a theory of imagination. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 14(1):15–23, March 1983. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170636.

Mittelstaedt:1983:WWL

[447] Peter Mittelstaedt. Wahrheit, Wirklichkeit und Logik in der Sprache der Physik. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 14(1):24–45, March 1983. CO-DEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170637.

Pfeiffer:1983:RVH

[448] K. Ludwig Pfeiffer. Rettung oder Verabschiedung der Hermeneutik? Funktionsgeschichte und Wirkungspotential neuerer hermeneutischer Denkfiguren. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 14(1):46–67, March 1983. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170638.

Scheibe:1983:TTS

[449] Erhard Scheibe. Two types of successor relations between theories. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 14(1):68–80, March 1983. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170639.

Rehder:1983:HVS

[450] Wulf Rehder. Hermeneutics versus Stupidities of All Sorts: A Review-Discussion of R. Rorty’s ‘Philosophy and the Mirror of Nature’. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 14(1):81–102, March 1983. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170640.

Danneberg:1983:IFDa

[451] Lutz Danneberg and Hans-Harald Müller. Der ‘intentionale Fehlschluß’: ein Dogma? Systematischer Forschungsbericht zur Kontroverse um eine intentionalistische Konzeption in den Textwissenschaften. Teil I. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 14(1):103–137, March 1983. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170641.

Engels:1983:EEB

[452] Eve-Marie Engels. Evolutionäre Erkenntnistheorie: ein biologischer Ausverkauf der Philosophie? Überlegungen zu Ortsbestimmung, Reichweite und Grenzen der Evolutionären Erkenntnistheorie im Anschluß an das 7. Internationale Wittgenstein-Symposium vom 22.–29. 8. 1982 in Kirchberg/Wechsel. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 14(1):138–166, March 1983. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170642.

Stuben:1983:RWI

[453] Peter E. Stüben. Rezension: Der Wissenschaftler und das Irrationale von Hans Peter Duerr. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 14(1):167–182, March 1983. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170643.

Coersmeier:1983:BG

[454] Birgit Coersmeier and Larry Steindler. [Bibliographie]. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 14(1):185–212, March 1983. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170644.

Anonymous:1983:FM

[455] Anonymous. Front matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 14(1):??, March 1983. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170634.

Anonymous:1983:VI

[456] Anonymous. Volume information. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 14(1):??, March 1983. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170633.

Bernier:1983:ORH

[457] Réjane Bernier. Origine et rôles de l’hypothèse en biologie. (French) [Origin and roles of the hypothesis in biology]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 14(2): 213–233, September 1983. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170646.

Fischer:1983:RHF

[458] Klaus Fischer. Rationale Heuristik: Die Funktion der Kritik im “Context of Discovery”. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 14(2):234–272, September 1983. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170647.

Grusser:1983:GBN

[459] O.-J. Grüsser. Über den Grad der Bewährung naturwissenschaftlicher Hypothesen. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 14(2):273–291, September 1983. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170648.

Holenstein:1983:SFG

[460] Elmar Holenstein. Zur Semantik der Funktionalanalyse. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 14(2):292–319, September 1983. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170649.

Keuth:1983:FVT

[461] Herbert Keuth. Fallibilismus versus transzendentalpragmatische Letztbegründung. (German) [Fallibilism versus transcendental pragmatic justification]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 14(2):320–337, September 1983. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170650.

VonLucadou:1983:PRS

[462] Walter Von Lucadou and Klaus Kornwachs. The problem of reductionism from a system theoretical viewpoint: How to link physics and psychology. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 14(2):338–349, September 1983. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170651.

Omer:1983:BDE

[463] I. A. Omer. Better deductive explanation? Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 14(2): 350–353, September 1983. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170652.

Stavenga:1983:FSP

[464] Gerben J. Stavenga. The fourth structure of physical reality. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 14(2):354–367, September 1983. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170653.

Carrier:1983:LBP

[465] Martin Carrier. Lakatos und Bohrs Programm: Entgegnung auf eine Kritik von Hans Radder. (German) [Lakatos and Bohr’s program: a response to a criticism of Hans Radder]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 14(2):368– 371, September 1983. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170654. See

[423].

Charpa:1983:NRI

[466] Ulrich Charpa. A note on rational inquiry in literary criticism. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 14(2):372–375, September 1983. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170655.

Danneberg:1983:IFDb

[467] Lutz Danneberg and Hans-Harald Müller. Der ‘intentionale Fehlschluß’: ein Dogma? Systematischer Forschungsbericht zur Kontroverse um eine intentionalistische Konzeption in den Textwissenschaften. Teil II. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 14(2):376–411, September 1983. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170656.

Anonymous:1983:EBG

[468] Anonymous. Eingegangene Bücher. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 14(2): 412–414, September 1983. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170657.

Drechsler:1984:GME

[469] Wolfgang Drechsler. Geometrie und Materie: Ist Einsteins Vision übertragbar auf die Elementarteilchenphysik?. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(1):1–21, March 1984. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170661.

Schmidt:1984:REL

[470] Klaus J. Schmidt. Ein Rechenverfahren für die elementare Logik II. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(1):22–33, March 1984. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170662.

Tietzel:1984:LMK

[471] Manfred Tietzel. L’homme machine: Künstliche Menschen in Philosophie, Mechanik und Literatur, betrachtet aus der Sicht der Wissenschaftstheorie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(1):34–71, March 1984. CO-DEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170663.

Welding:1984:SBW

[472] Steen Olaf Welding. Die Struktur der Begründung wissenschaftlicher Prognosen. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(1):72–91, March 1984. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170664.

Hovelmann:1984:SBE

[473] Gerd H. Hövelmann. Sprachkritische Bemerkungen zur evolutionären Erkenntnistheorie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(1):92–121, March 1984. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170665.

Janich:1984:IMH

[474] Peter Janich. Irrtum oder Methode? Zu den handlungstheoretischen Grundlagen der Protophysik in Auseinandersetzung mit der Kritik A. Kamlahs. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(1):122–141, March 1984. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170666.

Tetens:1984:PNG

[475] Holm Tetens. Physik am normativen Gängelband? Über das Verhältnis der Protophysik zur empirischen Physik. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(1):142–160, March 1984. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170667.

Heyer:1984:LWL

[476] Gerhard Heyer. Eine linguistische Wende in der Logik? Bericht über den 7. Internationalen Kongreß für Logik, Methodologie und Wissenschaftstheorie vom 11.–16. Juli 1983 in Salzburg. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(1):161–169, March 1984. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170668.

Hogrebe:1984:RSS

[477] Wolfram Hogrebe. Rezension: Sprache und Spracherlernung unter mathematisch-biologischer Perspektive, Berlin/New York 1981 von Volker Beeh. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(1):170–172, March 1984. CO-DEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170669.

Loose:1984:BRC

[478] Andreas Loose. Book review: Cosmology, Physics, and Philosophy by Benjamin Gal-Or. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(1):173–174, March 1984. CO-DEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170670.

Koppelberg:1984:BRW

[479] Dirk Koppelberg. Book review: Willard Van Orman Quine by Henri Lauener. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(1):174–177, March 1984. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170671.

Stranzinger:1984:RPW

[480] Rudolf Stranzinger. Rezension: Philosophie als Wissenschaft/Essays in Scientific Philosophy von Edgar Morscher, Otto Neumaier, Gerhard Zecha. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(1):177–178, March 1984. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170672.

Schmid:1984:RVG

[481] Michael Schmid. Rezension: Voraussetzungen und Grenzen der Wissenschaft von Gerard Radnitzky, Gunnar Andersson. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(1):179–184, March 1984. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170673.

Buttemeyer:1984:RNR

[482] Wilhelm Büttemeyer. Rezension: La nuova ragione: Scienza e cultura nella società contemporanea von Paolo Rossi. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(1):184–187, March 1984. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170674.

Coersmeier:1984:BGa

[483] Birgit Coersmeier and Larry Steindler. [Bibliographie]. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(1):188–196, March 1984. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170675.

Anonymous:1984:BMa

[484] Anonymous. Back matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(1):??, March 1984. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170676.

Anonymous:1984:FM

[485] Anonymous. Front matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(1):??, March 1984. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170660.

Anonymous:1984:VI

[486] Anonymous. Volume information. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(1):??, March 1984. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170659.

Hoyer:1984:LBG

[487] Ulrich Hoyer. Ludwig Boltzmann und das Grundlagenproblem der Quantentheorie. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(2):201–210, September 1984. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170678.

Loh:1984:VWI

[488] Werner Loh. Vorurteile und Wahn im logisch-mathematischen Grundlagenstreit und Probleme empirischer Begründung. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(2):211–231, September 1984. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170679.

Muller:1984:JVA

[489] Klaus F. O. Müller. Die jüngste Version einer älteren Weltanschauung: Eine Kritik an Paul Feyerabend. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(2): 232–260, September 1984. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170680.

Pearce:1984:RTI

[490] David Pearce. Research traditions, incommensurability and scientific progress. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(2):261–271, September 1984. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170681.

VanThemaat:1984:HDR

[491] W. A. Verloren Van Themaat. Hindsight and the definition of research success. illustrated by the history of planetary discovery. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(2):272–277, September 1984. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170682.

Weimann:1984:FNW

[492] Joachim Weimann. Die Funktion der normalen Wissenschaft in den Wirtschaftswissenschaften. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(2):278– 298, September 1984. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170683.

Zandvoort:1984:LNF

[493] Henk Zandvoort. Lakatos and Nagel: A Fruitful Confrontation. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(2):299–307, September 1984. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170684.

Busshoff:1984:WPW

[494] Heinrich Busshoff. Wissenschaftstheorie der Politischen Wissenschaft: Eine Antwort an Steinvorth. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(2):308– 329, September 1984. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170685.

Dieks:1984:RLC

[495] Dennis Dieks. The “reality” of the Lorentz contraction. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(2):330–342, September 1984. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170686.

Knight:1984:HSB

[496] David M. Knight. The history of science in Britain: A personal view. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(2):343–353, September 1984. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170687.

Jaenecke:1984:BRM

[497] Peter Jaenecke. Book review: Measurement. Its Concepts, Theories and Problems by Karel Berka. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(2):354–363, September 1984. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170688.

Schroder:1984:RTG

[498] Wilfried Schröder. Rezension: Theorie der Geowissenschaft von Wolf von Engelhardt, Jörg Zimmermann. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(2): 363–365, September 1984. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170689.

Rapp:1984:RBW

[499] Friedrich Rapp. Rezension: Die Bedeutung von ‘wahr’ und ‘Wahrheit’: Analysen zum Wahrheitsbegriff und zu einigen neueren Wahrheitstheorien von Winfried Franzen. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(2):365– 369, September 1984. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170690.

Rapp:1984:BRG

[500] Friedrich Rapp. Book review: A Guidebook for Technology Assessment and Impact Analysis by Alan L. Porter, Frederick A. Rossini, Stanley R. Carpenter, A. T. Roper, Ronal W. Larson, Jeffrey S. Tiller; A Guide to the Culture of Science, Technology, and Medicine by Paul T. Durbin. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(2):369–371, September 1984. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170691.

Coersmeier:1984:BGb

[501] Birgit Coersmeier and Larry Steindler. [Bibliographie]. (German) []. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(2):372–395, September 1984. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170692.

Anonymous:1984:BMb

[502] Anonymous. Back matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 15(2):??, September 1984. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170693.

Buksinski:1985:BHW

[503] Tadeusz Buksinski. Die Begründung des historischen Wissens. (German) [The grounding of historical knowledge]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(1):1–18, March 1985. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25170697.

Gadenne:1985:TBP

[504] Volker Gadenne. Theoretische Begriffe und die Prüfbarkeit von Theorien. (German) [Theoretical concepts and the testability of theories]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(1):19–24, March 1985. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170698.

Haga:1985:PSR

[505] Ånund Haga. Phenomenology and self-reflection. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(1):25–46, March 1985. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170699.

Idan:1985:TEP

[506] Asher Idan and Aharon Kantorovich. Towards an evolutionary pragmatics of science. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(1):47–66, March 1985. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170700.

Kramer:1985:FRP

[507] Hans Krämer. Funktions- und Reflexionsmöglichkeiten der Philosophiehistorie: Vorschläge zu ihrer wissenschaftstheoretischen Ortsbestimmung. (German) [Functional and reflective possibilities of the history of philosophy: Suggestions for their epistemological positioning]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(1):67–95, March 1985. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170701.

Widdershoven:1985:HSU

[508] Guy A. M. Widdershoven. Handlung und Struktur: Überlegungen zu einer handlungstheoretischen Gesellschaftstheorie. (German) [Action and structure: Reflections on an action-theoretical theory of society]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(1):96–112, March 1985. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170702.

Engels:1985:WLE

[509] Eve-Marie Engels. Was leistet die evolutionäre Erkenntnistheorie? Eine Kritik und Würdigung. (German) [what can evolutionary epistemology do? A review and assessment]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(1):113–146, March 1985. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170703.

Arens:1985:NSA

[510] Bernhard Arens. Die Non-standard Analysis: Eine Rehabilitierung des Unendlichkleinen in den Grundlagen der Mathematik. (German) [The non-standard analysis: a rehabilitation of the infinitely small in the foundations of mathematics]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(1):147–150, March 1985. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170704.

Gregory:1985:HIS

[511] Frederick Gregory. The historical investigation of science in North America. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(1):151–166, March 1985. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170705.

Wettersten:1985:BRN

[512] John R. Wettersten. Book review: New Methods for the Study of Man. Review Essay. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(1):167–176, March 1985. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170706.

Zittlau:1985:RWB

[513] Dieter Zittlau. Rezension: Wahrscheinlichkeitstheorie von Bruno de Finetti. (German) [Review: Probability Theory by Bruno de Finetti]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(1):176–177, March 1985. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170707.

Kornbichler:1985:RHV

[514] Thomas Kornbichler. Rezension: Historische Vernunft von Jörn Rüsen. (German) [Review: Historical Reason by Jörn Rüsen]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(1):177–182, March 1985. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170708.

Schafer:1985:RTW

[515] Lothar Schäfer. Rezension: Theorienwandel in der Wissenschaftsgeschichte. Chemie im 18. Jahrhundert von Elisabeth Ströker. (German) [Review: Change of theories in the history of science. Chemistry in the 18th Century by Elisabeth Ströker]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(1):182– 187, March 1985. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170709.

Dodel:1985:BGa

[516] Richard Dodel, Lothar Karschny, and Ricarda Van De Groote Poort. [Bibliographie]. (German) [[Bibliography]]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(1):188– 200, March 1985. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170710.

Anonymous:1985:FM

[517] Anonymous. Front matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(1):??, March 1985. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170696.

Anonymous:1985:VI

[518] Anonymous. Volume information. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(1):??, March 1985. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170695.

Carrier:1985:RBI

[519] Martin Carrier. Rudjer Boscovich und die induktive Logik. (German) [Rudjer Boscovich and inductive logic]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(2):201– 212, September 1985. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170712.

Diehl:1985:IDW

[520] Bardo Diehl. Intersystemarer Dialog in Wissenschaftstheorie und - geschichte. (German) [Intersystem dialogue in science theory and history]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(2):213–228, September 1985. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170713.

Engel-Tiercelin:1985:LPM

[521] Claudine Engel-Tiercelin. Logique, psychologie et métaphysique: les fondements du pragmatisme selon C. S. Peirce. (French) [Logic, psychology and metaphysics: the basis of pragmatism according to C. S. Peirce].

Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(2):229–250, September 1985. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170714.

Gauthier:1985:LAM

[522] Yvon Gauthier. The logical analysis of mathematical physics: Remarks on renormalization procedures in quantum field theories. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(2):251–260, September 1985. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170715.

Gavroglu:1985:PTS

[523] Kostas Gavroglu. Popper’s tetradic schema, progressive research programs, and the case of parity violation in elementary particle physics 1953–1958. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(2):261–286, September 1985. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170716.

Hoche:1985:SHM

[524] Hans-Ulrich Hoche. Sein und Heißen: Modalprobleme der Identität. (German) [Being and being called: modal problems of identity]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(2):287–303, September 1985. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170717.

Stock:1985:BZW

[525] Wolfgang G. Stock. Die Bedeutung der Zitatenanalyse für die Wissenschaftsforschung. (German) [The importance of citation analysis for science studies]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(2):304–314, September 1985. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170718.

Strube:1985:AVG

[526] Werner Strube. Analyse des Verstehensbegriffs. (German) [Analysis of the concept of understanding]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(2):315–333, September 1985. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170719.

Sundaram:1985:OLS

[527] K. Sundaram. Objectivity and the language of science. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(2):334–340, September 1985. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170720.

Albert:1985:MTM

[528] Hans Albert. Münchhausen in transzendentaler Maskerade: Über einen neuen Versuch der Letztbegründung praktischer Sätze. (German) [Münchausen in transcendental masquerade: On a new attempt at the ultimate justification of practical propositions]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(2): 341–356, September 1985. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170721.

Kuhlmann:1985:RLV

[529] Wolfgang Kuhlmann. Reflexive Letztbegründung versus radikaler Fallibilismus: Eine Replik. (German) [Reflective ultimate reasoning versus radical fallibilism: a replica]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(2):357–374, September 1985. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170722.

Scherer:1985:BRR

[530] Bernd Michael Scherer. Book review: Rational Discourse and Poetic Communication. Methods of Linguistic, Literary, and Philosophical Analysis by Roland Posner. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(2):375–380, September 1985. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170723.

Dodel:1985:BGb

[531] Richard Dodel, Lothar Karschny, and Ricarda Van De Groote Poort. [Bibliographie]. (German) [Bibliography]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 16(2):381– 396, September 1985. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170724.

Hoyer:1986:GRG

[532] Ulrich Hoyer. Die Grundlagen der Relativitätstheorie. (German) [The basis of relativity]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(1):1–13, March 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170728.

Loeck:1986:SEC

[533] Gisela Loeck. Ist Simulation Erklärung? Cognitive Science: wissenschaftstheoretisch betrachtet. (German) [Is simulation explanation? Cognitive science: viewed from a scientific-theoretical point of view]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(1):14–39, March 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170729.

Maaseide:1986:PIR

[534] Atle Måseide. Perspectivity, intentionality, reflection: Sketch for a theory of the life-world foundation of philosophy. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(1): 40–58, March 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170730.

Schopman:1986:NCF

[535] Joop Schopman. Negative cross-fertilization. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(1): 59–67, March 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170731.

Stockler:1986:PMR

[536] Manfred Stöckler. Philosophen in der Mikrowelt: ratlos? Zum gegenwärtigen Stand des Grundlagenstreits in der Quantenmechanik. (German) [Philosophers in the micro world: at a loss? On the current status of the fundamental dispute in quantum mechanics]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(1):68–95, March 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170732.

Tetens:1986:QVP

[537] Holm Tetens. “Ist die Quantenphysik vollständig?” Ein Plädoyer, die Kopenhagener Deutung der Quantenmechanik nicht realistischontologisch zu interpretieren. (German) [“Is quantum physics complete?” A plea for the Copenhagen not realistic-ontologic interpretation of quantum mechanics]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(1):96–118, March 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170733.

VanThemaat:1986:HFS

[538] W. A. Verloren Van Themaat. In how far is science accumulative? Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(1):119–130, March 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170734.

Wandschneider:1986:IEA

[539] Dieter Wandschneider. Die Inkonsistenz empiristischer Argumentation im Zusammenhang mit dem Problem der Naturgesetzlichkeit. (German) [The inconsistency of empiricist reasoning related to the problem of natural law]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(1):131–142, March 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170735.

Ginev:1986:SMI

[540] Dimíter Ginev. The “Science — methodology” iterative cycle: an analysis of A. Polikarov’s methodological views. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(1): 143–153, March 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170736.

Moser:1986:SJC

[541] Paul K. Moser. On scientific justification by consensus. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(1):154–161, March 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170737.

Gil:1986:RPZ

[542] Thomas Gil. Rezension: Philosophie als Zivilisationskritik · D. Böhlers Anatomie der dialogischen Vernunft. D. Böhler, Rekonstruktive Pragmatik. Von der Bewußtseinsphilosophie zur Kommunikationsreflexion: Neubegründung der praktischen Wissenschaften und Philosophie. (German) [Review: Philosophy as a critique of civilization · D. Böhler’s anatomy of dialogical reason. D. Böhler, Reconstructive Pragmatics. From the Philosophy of Consciousness to Reflection on Communication: Refounding the Practical Sciences and Philosophy]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(1): 162–165, March 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170738.

Flacke:1986:RAL

[543] Michael Flacke. Rezension: Analytische Literaturwissenschaft von Peter Finke, Siegfried J. Schmidt. (German) [Review: Analytical Literature Studies by Peter Finke, Siegfried J. Schmidt]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(1): 165–172, March 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170739.

Dodel:1986:BG

[544] Richard Dodel, Lothar Karschny, and Ricarda Van De Groote Poort. [Bibliographie]. (German) [Bibliography]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(1):173– 196, March 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170740.

Anonymous:1986:EBGa

[545] Anonymous. Eingegangene Bücher. (German) [Books received]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(1):197–200, March 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170741.

Anonymous:1986:FM

[546] Anonymous. Front matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(1):??, March 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170727.

Anonymous:1986:VI

[547] Anonymous. Volume information. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(1):??, March 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170726.

Carrier:1986:WRR

[548] Martin Carrier. Wissenschaftsgeschichte, rationale Rekonstruktion und die Begründung von Methodologien. (German) [History of science, rational reconstruction and the justification of methodologies]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(2):201–228, September 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170743.

Charpa:1986:PFG

[549] Ulrich Charpa. Philologischer Fortschritt. (German) [Philological progress]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(2):229–255, September 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170744.

Martens:1986:REV

[550] Bernd Martens. Reproduzierbarkeit von Ergebnissen vs. heuristischer Gehalt wissenschaftlicher Konzepte: Ein Fallbeispiel aus der Ökologie und allgemeinen Systemtheorie. (German) [Reproducibility of results vs. heuristic content of scientific concepts: a case study from ecology and general systems theory]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(2):256–264, September 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170745.

Mittelstaedt:1986:BIS

[551] Peter Mittelstaedt. Benennung und Identität in der Sprache der Physik. (German) [Naming and identity in the language of physics]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(2):265–294, September 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170746.

Nierlich:1986:WHE

[552] Edmund Nierlich. Was heißt es, eine empirisch-wissenschaftliche Theorie zu konstruieren?. (German) [What does it mean to construct an empiricalscientific theory?]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(2):295–314, September 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170747.

Tietzel:1986:IEG

[553] Manfred Tietzel. Idealisierte Erklärungen. (German) [Idealized explanations]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(2):315–321, September 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170748.

Weinert:1986:NGG

[554] Friedel Weinert. Nominalismus und Gesellschaft. (German) [Nominalism and society]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(2):322–345, September 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170749.

Schopman:1986:AIP

[555] Joop Schopman. Artifical intelligence and its paradigm. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(2):346–352, September 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170750.

Trimi-Kaleri:1986:EGL

[556] Katharina Trimi-Kaleri. Einige Gedanken zur Logik und ihrer Vermittlung aus gegebenem Anlaß. (German) [Some thoughts on logic and its mediation for the cause]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(2):353–360, September 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170751.

Wendel:1986:GLU

[557] Hans J. Wendel. Gibt es logisch unvereinbare aber dennoch empirisch äquivalente Gesamttheorien über die Welt?. (German) [Are there logically incompatible but nevertheless empirically equivalent overall theories about the world?]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(2):361–379, September 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170752.

Bernier:1986:BRF

[558] Réjane Bernier. Book review: Final and Efficient Causes. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(2):380–383, September 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170753.

Geldsetzer:1986:RUZ

[559] Lutz Geldsetzer. Rezension: Die Überwindung des Zufalls. Kritische Betrachtungen zu Leonard Nelsons Begründung der Ethik als Wissenschaft von Grete Henry-Hermann. (German) [Review: Overcoming Chance. Critical Reflections on Leonard Nelson’s Foundation of Ethics as Science by Grete Henry-Hermann]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(2):384–386, September 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170754.

Birnbacher:1986:RAP

[560] Dieter Birnbacher. Rezension: Analytische Philosophie von Hans-Ulrich Hoche, Werner Strube. (German) [Review: Analytical Philosophy by Hans-Ulrich Hoche, Werner Strube]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(2):387– 390, September 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170755.

Anonymous:1986:EBGb

[561] Anonymous. Eingegangene Bücher. (German) [Books received]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(2):391–395, September 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170756.

Anonymous:1986:BM

[562] Anonymous. Back matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 17(2):??, September 1986. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170757.

Geldsetzer:1987:AD

[563] Lutz Geldsetzer and Gert König. Alwin Diemer (1920–1986). Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1–2):1–21, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170761.

Agassi:1987:WHP

[564] Joseph Agassi. Whatever happened to the positivist theory of meaning. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1–2):22–29, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170762.

Agazzi:1987:STA

[565] Evandro Agazzi. A systems-theoretic approach to the problem of the responsibility of science. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1–2):30–49, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170763.

Bartels:1987:KVE

[566] Andreas Bartels. Kausalität ohne Vorhersagbarkeit: Eine These des Empirismus im Konflikt mit der Allgemeinen Relativitätstheorie. (German) [Causality without predictability: a thesis of empiricism in conflict with general relativity]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1–2):50–60, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170764.

Bayertz:1987:WEE

[567] Kurt Bayertz. Wissenschaftsentwicklung als Evolution? Evolutionäre Konzeptionen wissenschaftlichen Wandels bei Ernst Mach, Karl Popper und Stephen Toulmin. (German) [Science development as evolution? Evolutionary conceptions of scientific change in Ernst Mach, Karl Popper and Stephen Toulmin]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1–2):61–91, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170765.

Buchel:1987:KEG

[568] Wolfgang Büchel. Kausalität und Entropie. (German) [Causality and entropy]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1–2):92–95, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170766.

Druwe:1987:SIN

[569] Ulrich Druwe. The structural identity of the natural and social sciences. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1–2):96–109, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170767.

Freundlieb:1987:HS

[570] Dieter Freundlieb. Hermeneutics and semantics. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1– 2):110–133, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170768.

Herzog:1987:WWV

[571] Walter Herzog. Wissenschaft und Wissenschaftstheorie: Versuch einer Neubestimmung ihres Verhältnisses am Beispiel der Pädagogik. (German)

[Science and philosophy of science: attempt to redefine their relationship using the example of pedagogy]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1–2):134–164, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25170769.

Liegener:1987:CVP

[572] Christoph Liegener and Giuseppe Del Re. Chemistry vs. physics, the reduction myth, and the unity of science. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1–2): 165–174, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170770.

Lugg:1987:PPR

[573] Andrew Lugg. ‘The priority of paradigms’ revisited. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1–2):175–182, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170771.

Mendonca:1987:MKP

[574] W. P. Mendonça. Zur Möglichkeit kognitiver Psychologie aus Wittgensteinscher Sicht. (German) [On the possibility of cognitive psychology from a Wittgensteinian point of view]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1–2): 183–203, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170772.

Ohrstrom:1987:GMR

[575] Peter Øhrstrøm. Geochronology as a metaphysical research programme. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1–2):204–214, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170773.

Stekeler-Weithofer:1987:UAS

[576] Pirmin Stekeler-Weithofer. Sind die Urteile der Arithmetik synthetisch a priori? Zur sprachanalytischen Interpretation einer vernunftkritischen Überlegung. (German) [Are the judgments of arithmetic synthetic a priori? On the language-analytical interpretation of a reason-critical consideration]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1–2):215–238, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170774.

Strohmeyer:1987:TGT

[577] Ingeborg Strohmeyer. Tragweite und Grenze der Transzendentalphilosophie zur Grundlegung der Quantenphysik. (German) [Scope and limit of transcendental philosophy to the foundation of quantum physics]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1–2):239–275, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170775.

Bohme:1987:PEE

[578] Gernot Böhme. Professor Eberleins Entfinalisierung der Wissenschaftsphilosophie. (German) [Professor Eberlein’s de-finalization of the philosophy of science]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1–2):276–284, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170776.

Katthage:1987:EKG

[579] Karl-Heinrich Katthage. Die Eindeutigkeit der konstruktiven Geometrie. (German) [The uniqueness of constructive geometry]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1–2):285–295, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170777.

Rings:1987:SPT

[580] Hanspeter Rings. Das strukturalistische Problem der theoretischen Begriffe und seine Lösung: Kritische Bemerkungen. (German) [The structuralist problem of theoretical concepts and its solution: Critical remarks]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1–2):296–312, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170778.

Struve:1987:PET

[581] Horst Struve. Die Prüfbarkeit empirischer Theorien. (German) [The testability of empirical theories]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1–2):313–315, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170779.

Burrichter:1987:TIS

[582] C. Burrichter, R. Inhetveen, and W. Ch. Zimmerli. Technikphilosophie im Systemvergleich. (German) [Technology philosophy in system comparison]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1–2):316–321, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170780.

Wettersten:1987:BRN

[583] John Wettersten. Book review: New Problems of Technology; A Comprehensive Analysis. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1–2):322–331, ???? 1987. CO-DEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170781.

Hoven:1987:RWM

[584] Arno Hoven. Rezension: Die Wahrheit des Mythos von Kurt Hübner. (German) [Review: The truth of the myth by Kurt Hübner]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1–2):331–336, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170782.

Kornbichler:1987:RIK

[585] Thomas Kornbichler. Rezensionen: Die Idee einer Kritik der historischen Vernunft von Hans-Ulrich Lessing; Texte zur Kritik der historischen Vernunft von Wilhelm Dilthey. (German) [Reviews: The Idea of a Critique of Historical Reason by Hans-Ulrich Lessing; Texts on the Critique of Historical Reason by Wilhelm Dilthey]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1–2): 336–338, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170783.

Regelmann:1987:RIE

[586] Johann Peter Regelmann. Rezension: Isis entschleiert? Zum Verhältnis von biologischer und kultureller Evolution. (German) [Review: Isis unveiled? On the relationship between biological and cultural evolution]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1–2):338–348, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170784.

Nobis:1987:RPL

[587] Helmut Nobis. Rezension: Positionen der Literaturwissenschaft von David E. Wellbery. (German) [Review: Positions in Literary Studies by David E. Wellbery]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1–2):348–354, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170785.

Anonymous:1987:BSA

[588] Anonymous. Bibliographie der Schriften von Alwin Diemer. (German) [Bibliography of the writings of Alwin Diemer]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1– 2):355–359, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170786.

Dodel:1987:BG

[589] Richard Dodel, Arno Hoven, and Larry Steindler. [Bibliographie]. (German) [Bibliography]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1–2):360–394, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170787.

Anonymous:1987:BM

[590] Anonymous. Back matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1–2):??, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170788.

Anonymous:1987:FM

[591] Anonymous. Front matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1–2):??, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170760.

Anonymous:1987:VI

[592] Anonymous. Volume information. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 18(1–2):??, ???? 1987. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170759.

Addis:1988:DMS

[593] Laird Addis. Dispositional mental states: Chomsky and Freud. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(1):1–17, March 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170792.

Ginev:1988:SMS

[594] Dimiter Ginev and Asarja Polikarov. The scientification of methodology of science. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(1):18–27, March 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170793.

Hoyer:1988:TLG

[595] Ulrich Hoyer. Theorie der Lorentztransformationen. (German) [Theory of Lorentz transformations]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(1):28–36, March 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170794.

Speller:1988:AMT

[596] Jules Speller. Ein Argumentationsspiel um das Münchhausen-Trilemma. (German) [An argument game about the Münchausen trilemma]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(1):37–61, March 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170795.

Sukale:1988:LPG

[597] Michael Sukale. Logik und Psychologismus. (German) [Logic and psychologism]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(1):62–85, March 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170796.

Wandschneider:1988:KPR

[598] Dieter Wandschneider. Kants Problem der Realisierungsbedingungen organischer Zweckmäßigkeit und seine systemtheoretische Auflösung. (German) [Kant’s problem of the realization conditions of organic expediency and its system-theoretical solution]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(1):86– 102, March 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170797.

Agassi:1988:IG

[599] Joseph Agassi. Ixmann and the Gavagai. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(1):103– 116, March 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170798.

Quine:1988:CAR

[600] Willard Van Orman Quine. A comment on Agassi’s remarks. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(1):117–118, March 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170799.

Gethmann:1988:SAW

[601] Carl Friedrich Gethmann. Stolpersteine auf dem argumentativen Weg zur Logik: Bemerkungen zu einem Entwurf von P. Zahn. (German) [Stumbling blocks on the argumentative path to logic: Remarks on a draft by P. Zahn]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(1):119–123, March 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170800.

Schurz:1988:WPB

[602] Gerhard Schurz and Georg J. W. Dorn. Why Popper’s basic statements are not falsifiable some paradoxes in Popper’s “Logic of scientific discovery”. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(1):124–143, March 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170801.

Geldsetzer:1988:ECH

[603] Lutz Geldsetzer. Erster chinesischer Hermeneutikkongreß September 1987 in Shen Zhen. (German) [First Chinese Hermeneutics Congress September 1987 in Shen Zhen]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(1):144–147, March 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170802.

Furlong:1988:BRC

[604] John J. Furlong. Book review: Carter’s Cartesian cognitivism: Richard B. Carter, Descartes’ Medical Philosophy: The Organic Solution to the Mind-Body Problem. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(1):148–153, March 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170803.

Schopman:1988:BRE

[605] Joop Schopman. Book review: The Epistemological Interest of Artificial Intelligence. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(1):153–156, March 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170804.

Kitchener:1988:BRP

[606] Richard F. Kitchener. Book review: Psychogenèse et Histoire des Sciences par Jean Piaget, Rolando Garcia. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(1):157– 165, March 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170805.

Seibt:1988:RGW

[607] Johanna Seibt. Rezension: Die Grenzen der Wissenschaft von Nicholas Rescher. (German) [Review: The Frontiers of Science by Nicholas Rescher]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(1):165–170, March 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170806.

Coersmeier:1988:BG

[608] Birgit Coersmeier, Richard Dodel, Arno Hoven, and Larry Steindler. [Bibliographie]. (German) [Bibliography]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(1):171– 203, March 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170807.

Anonymous:1988:BMa

[609] Anonymous. Back matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(1):??, March 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170808.

Anonymous:1988:FM

[610] Anonymous. Front matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(1):??, March 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170791.

Anonymous:1988:VI

[611] Anonymous. Volume information. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(1):??, March 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170790.

Carrier:1988:NFD

[612] Martin Carrier. On novel facts: a discussion of criteria for non-ad-hoc-ness in the methodology of scientific research programmes. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(2):205–231, September 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170810.

Grusser:1988:ENK

[613] Otto-Joachim Grüsser. Ein Erkenntnismodell des Nikolaus von Kues und der Grad der Bewährung einer wissenschaftlichen Hypothese. (German) [An epistemological model of Nikolaus von Kues and the degree of validation of a scientific hypothesis]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(2):232–238, September 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (elec-

tronic). URL http://www.jstor.org/stable/25170811.

Kertesz:1988:BPE

[614] András Kertész. Zur Bewertung der pragmatischen Erklärungsmodelle. (German) [To evaluate the pragmatic explanatory models]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(2):239–251, September 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170812.

Lembeck:1988:WRI

[615] Karl-Heinz Lembeck. Wahrheitsähnlichkeit als Regulativ der Intentionalität? Ein Beispiel zum Verhältnis von Phänomenologie und moderner Wissenschaftstheorie. (German) [Truth-likeness as a regulator of intentionality? An example of the relationship between phenomenology and modern philosophy of science]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(2):252–265, September 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170813.

Loh:1988:UNW

[616] Werner Loh. Zur Überwindung neuzeitlicher Wissenschaftsauffassungen. (German) [To overcome modern scientific concepts]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(2):266–289, September 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170814.

Mainzer:1988:SSE

[617] Klaus Mainzer. Symmetrie und Symmetriebrechung: Zur Einheit und Vielheit in den modernen Naturwissenschaften. (German) [Symmetry and symmetry breaking: On unity and multiplicity in modern natural sciences]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(2):290–307, September 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170815.

Mittelstrass:1988:PWV

[618] Jürgen Mittelstrass. Die Philosophie der Wissenschaftstheorie: Über das Verhältnis von Wissenschaftstheorie, Wissenschaftsforschung und Wissenschaftsethik. (German) [The philosophy of the theory of science: on the relationship between the theory of science, the study of science and the ethics of science]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(2):308–327, September 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170816.

Spinner:1988:BIN

[619] Helmut F. Spinner. Die Besteigung des Informationsberges als neue Aufgabe der Philosophie im Verbund aller Wissenswissenschaften: Themen und Thesen zur philosophischen Bewältigung des Wandels der Wissensordnung infolge der informationstechnischen Entwicklung. (German) [Climbing the mountain of information as a new task for philosophy in the network of all knowledge sciences: Topics and theses on the philosophical management of the change in the knowledge system as a result of information technology development]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(2): 328–347, September 1988. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170817.

Hosle:1988:TGE

[620] Vittorio Hösle. Tragweite und Grenzen der evolutionären Erkenntnistheorie. (German) [Scope and limits of evolutionary epistemology]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(2):348–377, September 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170818.

Keuth:1988:FSG

[621] Herbert Keuth. Fehlbarkeit oder Sicherheit. (German) [Fallibility or security]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(2):378–390, September 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170819.

Ginev:1988:BRS

[622] Dimiter Ginev. Book review: The Search for a Methodology of Social Science by Stephen Turner. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(2):391–393, September 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170820.

Anonymous:1988:EBG

[623] Anonymous. Eingegangene Bücher. (German) [Books received]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(2):393–397, September 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170821.

Anonymous:1988:BMb

[624] Anonymous. Back matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 19(2):??, September 1988. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170822.

Dilworth:1989:NSL

[625] Craig Dilworth. On the nature of scientific laws and theories. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(1):1–17, March 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170826.

Gavroglu:1989:TTE

[626] Kostas Gavroglu. A taxonomy of theoretical and experimental tests. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(1):18–39, March 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170827.

Gregersen:1989:NTH

[627] Frans Gregersen and Simo Køppe. A normative theory of humanistic knowledge. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(1):40–53, March 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170828.

Kurthen:1989:III

[628] Martin Kurthen. Indeterminiertheit, Iterabilität und Intentionalität: Searle, Quine und Derrida zur Einheit der Bedeutung. (German) [Indeterminacy, iterability and intentionality: Searle, Quine and Derrida on the unity of meaning]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(1):54–86, March 1989. CO-DEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170829.

Lenk:1989:PWE

[629] Hans Lenk. Pragmatische Wende und “Erklärung” in der Wissenschaftstheorie. (German) [Pragmatic turn and “explanation” in philosophy of science]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(1):87–96, March 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170830.

Schopman:1989:TC

[630] Joop Schopman. The two cultures. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(1):97– 105, March 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170831.

Taschdjian:1989:SN

[631] Edgar Taschdjian. The sociotectonics of the noosphere. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(1):106–115, March 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170832.

VanThemaat:1989:UCB

[632] W. A. Verloren Van Themaat. Universal or culture-bound science? Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(1):116–123, March 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170833.

Albert:1989:HSO

[633] Hans Albert. Hösles Sprung in den objektiven Idealismus: Über die Verwirrungen eines ganz gewöhnlichen Genies. (German) [Hösle’s leap into objective idealism: On the confusions of an ordinary genius]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(1):124–131, March 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170834.

Beckermann:1989:BKM

[634] Ansgar Beckermann. Bedeutungsverstehen als Kennzeichen des Mentalen. (German) [Understanding meaning as a hallmark of the mental]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(1):132–145, March 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170835.

Gadenne:1989:ISA

[635] Volker Gadenne. Zur Interpretation der strukturalistischen Auffassung über die Prüfbarkeit von Theorien: Bemerkungen zum Beitrag von Horst Struve. (German) [On the interpretation of the structuralist view of the testability of theories: comments on Horst Struve’s contribution]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(1):146–147, March 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170836.

Bachmaier:1989:RPW

[636] Peter Bachmaier. Rezension: Die pragmatische Wende im ‘Widerstreit’?. (German) [Review: The pragmatic turn in ‘conflict’?]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(1):148–153, March 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170837.

Ginev:1989:RPK

[637] Dimiter Ginev. Rezension: Philosophieren mit Kant von Gernot Böhme. (German) [Review: Philosophy with Kant by Gernot Böhme]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(1):153–155, March 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170838.

Schopman:1989:BRW

[638] Joop Schopman. Book review: What Machines Can Do. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(1):156–161, March 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170839.

Schumacher:1989:BRI

[639] Ralph Schumacher. Book review: Individuality. An Essay on the Foundations of Metaphysics by Jorge J. E. Gracia. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(1): 162–166, March 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-

8587 (electronic). URL http://www.jstor.org/stable/25170840.

Schopman:1989:RBE

[640] Joop Schopman. Rezension: Der Baum der Erkenntnis. Die biologischen Wurzeln des menschlichen Erkennens von Humberto R. Maturana, Francisco J. Varela. (German) [Review: The Tree of Knowledge. The Biological Roots of Human Cognition by Humberto R. Maturana, Francisco J. Varela]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(1):166–169, March 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170841.

Buldt:1989:BG

[641] Bernd Buldt, Richard Dodel, and Arno Hoven. [Bibliographie]. (German) [Bibliography]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(1):170–197, March 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170842.

Anonymous:1989:BM

[642] Anonymous. Back matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(1):??, March 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170843.

Anonymous:1989:FM

[643] Anonymous. Front matter. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(1):??, March 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170825.

Anonymous:1989:VI

[644] Anonymous. Volume information. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(1):??, March 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170824.

Herausgeber:1989:E

[645] Die Herausgeber. Editorial. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(2):199–200, September 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170845.

Anapolitanos:1989:TTM

[646] D. A. Anapolitanos. Theories and their models. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(2): 201–211, September 1989. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170846.

Arageorgis:1989:DTS

[647] Aristidis Arageorgis and Aristides Baltas. Demarcating technology from science: Problems and problem solving in technology. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(2):212–229, September 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170847.

Becker:1989:FEU

[648] Wolfgang Becker. Freges Erläuterung des Urteils. (German) [Frege’s explanation of judgment]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(2):230–248, September 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170848.

Gavroglu:1989:QMM

[649] Kostas Gavroglu and Yorgos Goudaroulis. Quantum mechanics and macroscopic quantum phenomena: The case of superconductivity and superfluidity. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(2):249–275, September 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170849.

Geiger:1989:TEA

[650] Gebhard Geiger. Technik und Erkenntnis: Zur allgemeinen methodologischen und erkenntnistheoretischen Bedeutung der technischen Wissenschaften. (German) [Technology and knowledge: On the general methodological and epistemological significance of the technical sciences]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(2):276–286, September 1989. CODEN ZAWTA2.

ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170850.

Hoyer:1989:VGG

[651] Ulrich Hoyer. Verallgemeinerte Gravitationstheorie. (German) [Generalized theory of gravity]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(2):287–302, September 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170851.

Mendonca:1989:ASS

[652] W. P. Mendonça. Abstrakte Sinnesphysiologie als spekulative Philosophie. (German) [Abstract sensory physiology as speculative philosophy]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(2):303–316, September 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170852.

Menne:1989:KUI

[653] Albert Menne. Die Kantische Urteilstafel im Lichte der Logikgeschichte und der modernen Logik. (German) [The Kantian table of judgments in the light of the history of logic and modern logic]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(2):317–324, September 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/stable/

25170853.

Struve:1989:EGG

[654] Horst Struve. Empirische Geometrie. (German) [Empirical geometry]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(2):325–339, September 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170854.

VanThemaat:1989:OCM

[655] W. A. Verloren Van Themaat. The own character of mathematics discussed with consideration of the proof of the four-color theorem. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(2):340–350, September 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170855.

Keita:1989:USF

[656] Lansana Keita. Are universal statements falsifiable? A critique of Popper’s falsifiability criterion. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(2):351–366, September 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic).

URL http://www.jstor.org/stable/25170856.

Koppelberg:1989:QAT

[657] Dirk Koppelberg. Quine and the analytic tradition: a reply to Professor Agassi. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(2):367–372, September 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170857.

Stockler:1989:RPP

[658] Manfred Stöckler. Rezension: Philosophie und Physik der Raum-Zeit von Jürgen Audretsch, Klaus Mainzer. (German) [Review: Philosophy and Physics of Space-Time by Jürgen Audretsch, Klaus Mainzer]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(2):373–376, September 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.org/

stable/25170858.

Sobolewski:1989:RGP

[659] Anette Sobolewski. Rezension: GenEthik. Probleme der Technisierung menschlicher Fortpflanzung von Kurt Bayertz. (German) [Review: GenEthik. Problems of mechanization of human reproduction by Kurt Bayertz]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(2):376–380, September 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170859.

Engel:1989:RKR

[660] Gerhard Engel. Rezension: Karl R. Popper von Lothar Schäfer. (German) [Review: Karl R. Popper by Lothar Schäfer]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(2): 380–386, September 1989. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170860.

Stranzinger:1989:REV

[661] Rudolf Stranzinger. Rezension: Erklären und Verstehen in der Wissenschaft von Gerhard Schurz. (German) [Review: Explaining and understanding in science by Gerhard Schurz]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(2): 387–389, September 1989. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170861.

Ginev:1989:RNO

[662] Dimitri Ginev. Rezension: Über Natur von Oswald Schwemmer. (German) [Review: On Nature by Oswald Schwemmer]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(2): 389–390, September 1989. CODEN ZAWTA2. ISSN 0044-2216 (print),

1572-8587 (electronic). URL http://www.jstor.org/stable/25170862.

Anonymous:1989:NBS

[663] Anonymous. Nachträge zur Bibliographie der Schriften von Alwin Diemer. (German) [Supplements to the bibliography of the writings of Alwin Diemer]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 20(2):391–394, September 1989. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170863.